### v2

# Dual Alignment Maximin Optimization for Offline Model-based RL

Abstract
Offline reinforcement learning agents face significant deployment challenges due to the synthetic-to-real distribution mismatch. While most prior research has focused on improving the fidelity of synthetic sampling and incorporating off-policy mechanisms, the directly integrated paradigm often fails to ensure consistent policy behavior in biased models and underlying environmental dynamics, which inherently arise from discrepancies between behavior and learning policies. In this paper, we first shift the focus from model reliability to policy discrepancies while optimizing for expected returns, and then self-consistently incorporate synthetic data, deriving a novel actor-critic paradigm, Dual Alignment Maximin Optimization (DAMO). It is a unified framework to ensure both model-environment policy consistency and synthetic and offline data compatibility. The inner minimization performs dual conservative value estimation, aligning policies and trajectories to avoid out-of-distribution states and actions, while the outer maximization ensures that policy improvements remain consistent with inner value estimates. Empirical evaluations demonstrate that DAMO effectively ensures model and policy alignments, achieving competitive performance across diverse benchmark tasks.
Machine Learning, Offline Reinforcement learning
1 Introduction
Offline reinforcement learning (RL) (Lange et al., 2012; Levine et al., 2020) aims to learn policies from a pre-collected dataset generated by a behavioral policy within the real environment. This paradigm helps avoid the safety risks and high costs associated with direct interactions, making RL applicable in real-life scenarios, such as healthcare decision-making support and autonomous driving (Emerson et al., 2023; Sinha et al., 2022; Mnih et al., 2015). Offline model-based RL (Yu et al., 2020; Kidambi et al., 2020; Sun et al., 2023) improves upon this by training a dynamics model and using it to generate synthetic data for policy training or planning. The introduction of dynamics models enhances the sample efficiency and allows the agent to answer counterfactual queries (Levine et al., 2020). However, despite strong performance in learned models, policies often degrade significantly when deployed in real environments due to mismatches between synthetic and real distributions.
Most prior works have focused on enhancing the reliability of synthetic samplings from learned models, and directly incorporating off-policy optimization methods, such as SAC (Haarnoja et al., 2018), to address this mismatch. The representative approach is MOPO (Yu et al., 2020), which quantifies model uncertainty by measuring the predicted variance of learned models and subsequently penalizes model-generated trajectories with high uncertainty. This methodology has inspired subsequent research like MOReL (Kidambi et al., 2020) and MOBILE (Sun et al., 2023), which adopt similar uncertainty-aware frameworks. These approaches help generate synthetic data that remains compatible with the offline dataset and construct conservative value functions to mitigate out-of-distribution (OOD) actions. However, creating a perfect learned model from limited offline datasets is impossible, and as illustrated in Fig. 1, simply combining these methods with off-policy mechanisms still fails to resolve the inherent discrepancies in policy behaviors between the learned model and the underlying environment, leading to inconsistent policy performance and leaving the distribution shift problem unsolved.
In this paper, we trace the inherent root of distribution shift challenges in offline RL to the discrepancies between behavior and learning policies in the underlying environmental dynamics. While optimizing for expected returns, we introduce a regularized policy optimization objective for offline RL, that constrains the visitation distribution discrepancies between the behavioral and the learning policies in real environments. We further self-consistently incorporate synthetic data into this objective, deriving a novel maximin optimization objective for offline model-based RL. Building on these objectives, we propose Dual Alignment Maximin Optimization (DAMO), a consistent actor-critic framework that ensures both coherent alignments between model and environment behavior and model-generated and offline data.
DAMO iteratively applies maximin optimization in two phases. In the inner minimization phase, equivalent to critic training, the focus is on the dual conservative value estimation. This value estimation implicitly harmonizes learning policy behaviors in both the model and the environment to avoid OOD states, while explicitly aligning synthetic and offline data to prevent OOD actions. In the outer maximization phase, corresponding to actor learning, policy improvements are guided by the inner value estimates through objective alignment. This preserves reliably conservative behavior and effectively mitigates OOD issues throughout both policy optimization and deployment. Additionally, DAMO incorporates classifiers to approximate the data alignment terms, facilitating practical implementation. Extensive experiments showcase that DAMO successfully mitigates OOD issues and achieves competitive performance across various benchmarks. Our contributions are summarized as the following:
  • We derive a consistent maximin object for offline model-based RL, ensuring both alignments between model and environment policy behaviors, and synthetic and offline data.
  • We develop DAMO, a novel actor-critic framework that employs maximin optimization to achieve the dual conservative value estimation and the consistent policy improvement.
  • We conduct extensive experiments to validate that DAMO effectively alleviates the distribution shift challenge, delivering superior performance across varying benchmark tasks.
(a) MOPO
(b) MOBILE
(c) DAMO
Figure 1: The distribution of transition pair $(s,a,s^{\prime})$ across three datasets (Offline Data, Real Data, and Synthetic Data) are shown for the hopper-medium-expert task, comparing two SOTA methods (MOPO and MOBILE) with our proposed DAMO framework. While MOPO and MOBILE align synthetic and offline data well (indicated by a high overlap), they fail to address OOD states in the real environment (circled by an ellipse drawn with a green dotted line). In contrast, DAMO effectively eliminates the OOD region.
2 Related Work
Offline Model-based RL trains a dynamics model using offline data to approximate environmental dynamics and performs conservative policy optimization (Lu et al., 2021) with model-generate data to mitigate OOD issues. MOPO (Yu et al., 2020) penalizes state-action pairs with high model uncertainty, effectively discouraging the agent from selecting OOD actions. MOReL (Kidambi et al., 2020) constructs a pessimistic Markov decision process to prevent the agent from entering OOD regions, ensuring safer and more reliable policy learning. COMBO (Yu et al., 2021) extends Conservative Q-learning (Kumar et al., 2020) to the model-based setting by regularizing the value function on OOD samples. RAMBO (Rigter et al., 2022) employs an adversarial training framework, optimizing both the policy and model jointly to ensure accurate transition predictions while maintaining robustness. MOBILE (Sun et al., 2023) incorporates penalties into value learning by utilizing uncertainty in Bellman function estimates derived from ensemble models. SAMBO (Luo et al., 2024a) imposes reward penalties and fosters exploration by incorporating model and policy shifts inferred through a probabilistic inference framework.
DICE-based Methods. Distribution Correction Estimation (DICE) is a technique that estimates stationary distribution ratios using duality theory. DICE has shown superior performance in evaluating discrepancies between distributions and has been widely applied in various RL domains, including off-policy evaluation (OPE) (Nachum et al., 2019a), offline imitation learning (IL) (Ma et al., 2022), and offline RL (Nachum et al., 2019b). In offline RL, DICE-based methods correct distribution shifts between offline data and environmental dynamics, formulating a tractable maximin optimization objective. For instance, AlgaeDICE (Nachum et al., 2019b) pioneers the application of DICE-based methods in offline RL, employing regularized dual objectives and Lagrangian techniques to address distribution shift challenges. OptiDICE (Lee et al., 2021) incorporates the Bellman flow constraint and directly estimates stationary distribution corrections for the optimal policy, eliminating the need for policy gradients. ODICE (Mao et al., 2024a) combines orthogonal-gradient updates with DICE to enforce state-action-level constraints. Unlike these model-free approaches, DAMO is, to the best of our knowledge, the first method to promote DICE insights in the model-based setting. While model-free methods typically handle discrepancies between two distributions, DAMO skillfully extends this concept by managing the differences across three distributions, using a divergence upper bound to ensure alignment.
Figure 2: Intuiative examples for demystifying OOD issues in offline model-based RL. The behavior policy $\pi_{\beta}$ primarily favors the top road ($a_{1}$) with a minor representation of the bottom road ($a_{3}$) and never selects the middle road ($a_{2}$). Consequently, the dynamics model lacks representation of medium road ($a_{2}$) and treats the $a_{3}$ as an OOD action. The policy $\pi$, trained on both offline and synthetic data, avoids $a_{3}$ when deployed in the real environment but struggles to evaluate $a_{2}$ due to its absence in $M$, resulting in occasionally getting into OOD state. In DAMO, $\pi$ is constrained by $\pi_{\beta}$, guiding the agent to consistently select $a_{1}$. At the bottom, we show that while data alignment ensures a similar distribution between offline and synthetic data, it fails to guarantee behavior consistency between $M$ and $T$.
3 Preliminaries
MDP. We consider the Markov decision process (MDP) defined by the six-element tuple $M=(\mathcal{S},\mathcal{A},T,r,\mu_{0},\gamma)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ denotes the action space, $T(s^{\prime}|s,a)$ is the environment transition dynamics, $r(s,a,s^{\prime})>0$ is the reward function, $\mu_{0}$ is the initial state distribution, and $\gamma\in(0,1)$ is the discounted factor. Given an MDP, the objective of RL is to find a policy $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})$ that maximizes the cumulative reward from the environment. This objective can be formally expressed as: $\pi^{\star}=\arg\max_{\pi}\mathbb{E}_{\mu_{0},\pi,T}[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t},s_{t+1})]$. In the derivation of DAMO, we utilize the dual form of the RL objective (Puterman, 2014), which is represented as follows:
$\pi^{\star}=\arg\max\limits_{\pi}\mathbb{E}_{\rho^{\pi}_{T}}[r(s,a,s^{\prime})].$
Here, $\rho^{\pi}_{T}(s,a,s^{\prime})$ is the transition occupancy measure, characterizing the distribution of transition pairs $(s,a,s^{\prime})$ induced by policy $\pi$ under the dynamics $T$, defined by:
$\rho_{T}^{\pi}(s,a,s^{\prime})=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}\mathbb{P}[s_{t}=s,a_{t}=a,\\ s_{t+1}=s^{\prime}|s_{0}\sim\mu_{0},a_{t}\sim\pi(\cdot|s_{t}),s_{t+1}\sim T(\cdot|s_{t},a_{t})].$
Here, $\mu_{0}$ denotes the initial state distribution. For simplicity, we use shorthand notation $\rho^{\pi}_{T}$ in the remainder of the article.
Offline Model-based RL. Offline model-based RL aims to find the optimal policy by leveraging a learned dynamics model $M$. Given the offline dataset $\mathcal{D}_{R}$, $M$ is typically trained using maximum likelihood estimation, formulated as $\min_{M}\mathbb{E}_{s,a,s^{\prime}\sim\mathcal{D}_{R}}[-\log M(s^{\prime}|s,a)]$. The training policy $\pi$ will undergo $k$-step rollout in learned dynamics model $M$ and the generated data will be added to a replay buffer $\mathcal{D}_{M}$. Then $\pi$ can be optimized using mini-batches of data sampled from $\mathcal{D}_{R}\cup\mathcal{D}_{M}$, where each datapoint is sampled from the offline dataset $\mathcal{D}_{R}$ with probability $f$ and the model dataset $\mathcal{D}_{M}$ with probability $1-f$. Throughout this work, we assume that the reward function $r(s,a,s^{\prime})$ is known, though it can also be treated as part of the model if unknown.
4 Demystify OOD in Offline Model-based RL
In this section, as illustrated in Fig. 2, we explore the out-of-distribution (OOD) issues in offline model-based RL. We focus on the importance of aligning both the model and offline data, as well as the model and environment policy behavior, which motivates the direction of this research.
4.1 OOD Issues in Offline RL
In offline RL, OOD problems (Mao et al., 2024b) arise when agents encounter states or select actions that fall outside the distribution of the offline data during training or testing.
OOD actions refer to actions that the behavior policy does not choose in specific states, rather than actions that are simply absent from the offline dataset. This implies that OOD actions are inherently state-dependent. Thus some studies use the term OOD state-action pairs to describe the state-action pair $(s,a)$ that does not appear in the offline dataset. Taking OOD actions during training can lead to severely inaccurate value estimation and significantly degraded policy performance in testing. Consequently, avoiding OOD actions has been a focus in previous offline RL research.
OOD states primarily arise in three main scenarios: 1) The learned policy executes unreliable OOD actions, leading to transitions into OOD states. 2) The initial state of the real environment lies outside the offline dataset. Additionally, stochastic dynamics could unexpectedly transition the agent into OOD states, even if it takes in-distribution (ID) actions in ID states. 3) Nonstationary dynamics, such as disturbances in real-world environments, can introduce unexpected changes in state distributions. While this factor is outside the scope of our offline model-based setting, it is an important consideration in real-world applications.
4.2 OOD Issues under Dynamics Model Integration
OOD issues in offline model-based RL are intrinsically linked to the mismatch between synthetic and real data distributions, compounded in model-based settings, where model inaccuracies introduce additional sources of error. More extensive discussions are provided in Appendix B.
OOD Actions. The model, $M$, is used to generate synthetic data for training, but it is essential to avoid taking OOD actions when rolling out policy in the model. This can trigger negative model exploitation (Levine et al., 2020), resulting in poor policy performance. Aligning the distributions of synthetic and offline data can effectively mitigate these issues as it ensures accurate value estimation.
OOD States. The performance of the learning policy during deployment is critically dependent on the alignment of transition pairs between the learned model and the real environment. Significant discrepancies between these transitions can cause performance degradation and lead to OOD states, highlighting the necessity of correcting transition biases during training. Even if the learned policy performs well in the model, it is important, during real-environment deployment, to avoid OOD states, for which we cannot accurately estimate the value function in the offline setting.
In summary, while optimizing the policy $\pi$ to maximize returns, we should constrain it in two directions: 1) aligning the synthetic data $(s,a,s^{\prime})_{M}^{\pi}$ with the offline data $(s,a,s^{\prime})_{T}^{\pi_{\beta}}$, and 2) ensuring that $\pi$ exhibits consistent behavior in both dynamics models and real environments.
5 Dual Alignment Maximin Optimization
Inspired by the analysis above, we propose Dual Alignment Maximin Optimization (DAMO), a unified approach to mitigate both OOD actions and states. We further explore the distinct roles of the inner minimization and the outer maximization in addressing the distribution shift challenge.
5.1 Maximin Objective for Unified Shift Mitigation
We start from the root cause of distribution shifts, i.e., the discrepancy between learned and behavior policy behaviors in the underlying environment. To tackle this, we introduce a regularized objective that constrains the difference between these behaviors while optimizing for expected rewards:
$\max\limits_{\pi}\mathbb{E}_{\rho_{T}^{\pi}}[r(s,a,s^{\prime})]-\alpha D_{KL}(\rho_{T}^{\pi}\|\rho_{T}^{\pi_{\beta}}),$
where $\alpha$ is a regularization hyperparameter. Next, we self-consistently incorporate synthetic data through a divergence upper bound, deriving the following surrogate objective:
$\max\limits_{\pi}\mathbb{E}_{\rho_{T}^{\pi}}[r(s,a,s^{\prime})-\alpha\log\dfrac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}]-\alpha D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi}).$ (1)
This surrogate objective integrates two essential elements: a data alignment regularizer $\log(\rho_{M}^{\pi}/\rho_{T}^{\pi_{\beta}})$ to ensure compatibility between synthetic and offline data, and a behavior alignment regularizer $D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})$ that maintains consistent policy performance across models and real environments.
While this objective (1) theoretically addresses the synthetic-to-real distribution mismatch (as shown later in Sec. 5.3), it is impractical to directly compute due to the inaccessibility of $\rho_{T}^{\pi}$. To make this objective tractable, we apply the convex conjugate theorem into the dual formulation of reinforcement learning, reformulating it as a maximin optimization problem. This reformulation forms the core of our proposed DAMO, a unified maximin framework for mitigating distribution shifts in offline model-based RL.
Theorem 5.1.
(1) is equivalent to the following problem:
$\max\limits_{\pi}\min\limits_{\tilde{Q}(s,a)}(1-\gamma)\mathrm{E}_{s\sim\mu_{0},a\sim\pi}[\tilde{Q}(s,a)]\\ +\alpha\mathrm{E}_{\rho_{M}^{\pi}}[f_{\star}(\Phi(s,a,s^{\prime})/\alpha)],$ (2)
where $f_{\star}$ is the conjugate function of the convex function $f$, and $\Phi(s,a,s^{\prime})=\log r(s,a,s^{\prime})-\alpha\log\frac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}+\tau^{\pi}Q(s,a)-Q(s,a)$, with $\tau^{\pi}Q(s,a)=\gamma\sum_{a^{\prime}}Q(s^{\prime},a^{\prime})\pi(a^{\prime}|s^{\prime})$. Here, $\mu_{0}$ denotes the distribution of the initial state, and $\alpha$ is a hyperparameter to control the degree of conservatism.
5.2 Practical Implementation of DAMO
To implement the maximin optimization (2) practically, we employ a classifier to approximate the data alignment term $\log(\rho_{M}^{\pi}/\rho_{T}^{\pi_{\beta}})$. Specifically, we train a classifier $h(s,a,s^{\prime})$ using the following loss function to distinguish transitions sampled from offline data and those generated by the model:
$\min_{h}\dfrac{1}{|\mathcal{D}_{R}|}\sum_{(s,a,s^{\prime})\in\mathcal{D}_{R}}\log h(s,a,s^{\prime})\\ +\dfrac{1}{|\mathcal{D}_{M}|}\sum_{(s,a,s^{\prime})\in\mathcal{D}_{M}}[\log(1-h(s,a,s^{\prime}))].$ (3)
Here, $\mathcal{D}_{R}$ and $\mathcal{D}_{M}$ represent the offline and the model-generated data, respectively. Based on the learned classifier $h^{\star}(s,a,s^{\prime})$, the data alignment modification $\log(\rho_{M}^{\pi}/\rho_{T}^{\pi_{\beta}})$ can be computed using the following analytical expression:
$\log\dfrac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}=\log\dfrac{h^{\star}(s,a,s^{\prime})}{1-h^{\star}(s,a,s^{\prime})}.$ (4)
Note that (2) can be divided into two phases: an inner value estimation phase and an outer policy improvement phase, resembling the structure of the actor-critic framework. This structure is directly adapted to implement DAMO. The complete DAMO framework is outlined in Algorithm 1.
Algorithm 1 DAMO
0:  Offline dataset $\mathcal{D}_{R}$, dynamics model $M$, actor network $\pi_{\phi}$, critic network $Q_{\theta}$, classifier $h_{\psi}$.
1:  Train the probabilistic dynamics model $M$ on $\mathcal{D}_{R}$
2:  Initialize the model dataset buffer $\mathcal{D}_{M}\leftarrow\varnothing$
3:  for each epoch do
4:     Generate synthetic rollouts by model $M$
5:     Add transition data in these rollouts to $\mathcal{D}_{M}$
6:     Update classifier $h_{\psi}$ according to Eq. (3)
7:     Calculate $\log(\rho_{M}^{\pi}/\rho_{T}^{\pi_{\beta}})$ according to Eq. (4)
8:     Update critic $Q_{\theta}$ via minimizing Eq. (2)
9:     Update actor $\pi_{\phi}$ via maximizing Eq. (2)
10:  end for
5.3 Theoretical Insights for DAMO
The core structure of DAMO involves mapping the inner minimization in (2) to the estimation of a dual conservative value $\tilde{Q}(s,a)$, and the outer maximization to policy improvement based on this value. In this subsection, we present deeper theoretical insights into the workings of DAMO.
Dual Conservative Value Estimation. We first demonstrate the existence of both explicit and implicit reward penalties in the optimization objective (2), which contribute to the construction of the dual conservative value estimation.
Theorem 5.2.
The optimal solution $Q_{\pi}^{\star}$ for the inner minimization optimization in (2) satisfies the following equation:
$Q_{\pi}^{\star}=r-\alpha\log\dfrac{\rho^{\pi}_{M}}{\rho^{\pi_{\beta}}_{T}}-\alpha f^{\prime}(\dfrac{\rho^{\pi}_{T}}{\rho^{\pi}_{M}})+\tau^{\pi}Q_{\pi}^{\star}.$
This indicates that DAMO inherently introduces an explicit data alignment modification, $\log(\rho^{\pi}_{M}/\rho^{\pi{\beta}}_{T})$, and an implicit behavior alignment adjustment, $f^{\prime}(\rho^{\pi}_{T}/\rho^{\pi}_{M})$. Minimizing the inner objective effectively reshapes the reward function, $r(s,a,s^{\prime})$, into a refined reward, $\tilde{r}(s,a,s^{\prime})$, which represents the dual conservative value estimation:
$\tilde{r}=r-\alpha\log\frac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}-\alpha f^{\prime}(\dfrac{\rho^{\pi}_{T}}{\rho^{\pi}_{M}}).$
The data alignment modification $\log(\rho^{\pi}_{M}/\rho^{\pi{\beta}}_{T})$ penalizes transition $(s,a,s^{\prime})$ that exhibit discrepancies between synthetic and offline data, promoting alignment with in-distribution transitions. This helps ensure accurate value estimation for these transitions. Meanwhile, the behavior alignment adjustment $f^{\prime}(\rho^{\pi}_{T}/\rho^{\pi}_{M})$ focuses on that different dynamics impact the distribution of $(s,a,s^{\prime})$. By penalizing these discrepancies, we encourage the agent to select policies that exhibit similar performance in both the underlying environment and the dynamics model, thus mitigating the impact of OOD states and ensuring consistent policy behavior between dynamics models and real environments.
Consistent Policy Improvement. Unlike conventional offline model-based approaches that optimize policies purely by maximizing estimated values, DAMO ensures the consistency between value estimation and policy improvement. Specially, DAMO employs $\tilde{Q}(s,a)$ to minimize the objective (2) and utilizes $\pi$ to maximize it. Conventional approaches typically follow a bilevel optimization framework, while DAMO employs a maximin paradigm, ensuring consistency and preserving dual conservatism. As demonstrated in Sec. 6.1, this consistent structure maintains the conservatism of value estimation during policy improvement.
Once the inner minimization in (2) is solved, the outer maximization problem is equivalent to solving the surrogate objective (1), which provides a lower bound for the standard RL objective $\mathcal{J}(\pi)=\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t},s_{t+1})]$:
Theorem 5.3.
(1) is a lower bound of $\mathcal{J}(\pi)$ for all $\pi$:
$\mathcal{J}(\pi)\geq\mathbb{E}_{\rho_{T}^{\pi}}[r(s,a,s^{\prime})-\alpha\log\dfrac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}]-\alpha D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi}).$
This demonstrates that the equivalence between objective (2) and the standard reinforcement learning objective ensures that the learned policy performs well in the environment.
6 Experiments
In this section, we focus on the following key aspects: 1) The capability of DAMO to handle synthetic-to-real distribution mismatch, as well as the contributions of its components. 2) The comparative performance of DAMO against existing methods on standard offline RL benchmarks. 3) The significant impact of hyperparameter settings and implementation approaches on the overall performance of DAMO.
We delve into these aspects using the D4RL benchmark (Fu et al., 2020) on the MuJoCo simulator (Todorov et al., 2012). Our implementation is based on the OfflineRL-Kit library [^1] https://github.com/yihaosun1124/OfflineRL-Kit, a comprehensive and high-performance library for implementing offline reinforcement learning algorithms. The basic parameters of DAMO are consistent with the settings of this library. The details of each experiment and the specific hyperparameters configuration can be found in Appendix D.
6.1 Effectiveness of DAMO Paradigm
In this section, we present empirical evidence demonstrating the dual capabilities of DAMO: 1) effectively aligning model-generated data with offline data distributions, and 2) consistently maintaining policy stability between dynamics models and real environments. Our detailed ablation study thoroughly investigates two key components of DAMO: inner value estimation and outer policy improvement.
(a) The distribution of w/o er
(b) The distribution of w/o ir
(c) The distribution of DAMO
Figure 3: (a) (b) (c) illustrate the results of DAMO without the data alignment term (w/o er), DAMO without the behavior alignment term (w/o ir), DAMO. The left panels depict the distributions of state-action pair $(s,a)$ for offline and synthetic data. The right panels display the distributions of states for real and synthetic data.
(a) Normalized Score
(b) Estimated Value
Figure 4: Training process of DAMO, w/o ir, w/o er. (a) The data alignment term is the primary factor that influences policy improvement, while the absence of the behavior alignment term results in performance degradation in the final stages. (b) The lack of the data alignment term leads to the overestimation of the value.
6.1.1 Dual Conservative Value Estimation
In this part, we present an ablation study examining the individual contributions of the data alignment and behavior alignment terms. We systematically evaluate their impact by removing each term from DAMO and training policies on the hopper-medium-expert-v2 benchmark. To provide intuitive insights, we first visualize the state-action distributions of the final trained policies across three distinct dataset categories. These categories are defined as follows:
  • Offline Data: The transition pairs $(s,a,s^{\prime})$ are collected with behavior policy $\pi_{\beta}$ in dataset Hopper-medium-expert.
  • Synthetic Data: The transition pairs are generated by conducting training policy $\pi$ within the dynamics model $M$.
  • Real Data: The transition pairs are generated by conducting training policy $\pi$ within the real environment $T$.
Effectiveness of Data Alignment. To evaluate synthetic and offline data compatibility during training, we analyze the state-action pair distributions for both synthetic data and offline data under three distinct conditions: the complete DAMO framework, the configuration without the behavior alignment term (w/o ir), and the configuration without the data alignment term (w/o er). The visualization results are presented in the left column of Fig. 3. Building on these initial observations, we then examine the training dynamics of DAMO under three specific conditions to further investigate their significant impacts. The results are presented in Fig. 4. A comparative analysis between Fig. 3(a) and Fig. 3(c) reveals that the w/o er configuration fails to establish proper alignment between synthetic data and offline data, thereby inadequately mitigating OOD state-action pairs. The results in Fig. 3(b) further highlight that the data alignment term enables w/o ir configuration to align synthetic data with offline data. To clarify the detrimental effects of incompatibility between these two data sources, Fig. 4(b) illustrates that w/o er overestimates the value of OOD actions, leading to suboptimal training performance. This is corroborated by the poor evaluation scores of w/o er depicted in Fig. 4(a).
Effectiveness of Behavior Alignment. To assess the model-environment consistency during the testing phase, we follow a similar evaluation procedure. We visualize the state distributions of real environment and model-generated data under the same conditions. The visualization results are presented in the right column of Fig. 3. A comparative analysis between Fig. 3(b) and Fig. 3(c) reveals that while w/o ir successfully achieves synthetic-offline data compatibility and mitigates OOD actions, the absence of the behavior alignment term leads to incomplete mitigation of OOD states, manifesting policy inconsistency between dynamics model and real environment. Although the behavior alignment term imposes constraints, the results in Fig. 3(a) demonstrate that w/o er fails to effectively mitigate OOD states. The observed discrepancies in state distributions between real environment data and model-generated data under the w/o er highlight the critical importance of establishing robust compatibility between synthetic and offline data. This compatibility is essential for ensuring consistent policy performance across dynamics models and real environments. The influence is reflected in Fig.4(a), which shows performance degradation of w/o ir in the final stage.
6.1.2 Towards Consistent Policy Improvement
Table 1: The average value of OOD state-action pairs, $\infty$ for severe overestimation. The smallest values are bolded, which indicates the best mitigation of the overestimation problem.
halfcheetah hopper walker2d
Consistent Version 95.2 213.0 231.7
Inconsistent Version 101.4 219.4 $\infty$
$Real\ value$ 819.6 269.6 299.7
This section presents a statistical analysis of the value estimation performance of DAMO across multiple datasets, with comparisons against the true value function. Our results validate that the compatible policy improvement framework effectively maintains the conservatism of value estimation.
To investigate the significance of objective alignment in DAMO, we conduct a controlled comparison experiment by directly integrating DAMO with standard off-policy mechanisms (Inconsistent Version) while evaluating the critic’s estimation of state-action pair values and comparing it with the original DAMO (Consistent Version). We perform this analysis on the medium-expert dataset across three distinct environments: HalfCheetah, Hopper, and Walker2d. The obtained quantitative results are summarized in Table 1. The experimental findings reveal that the Consistent Version achieves conservative value estimations across all environments. Although the Inconsistent version maintains conservatism in the HalfCheetah and Hopper environments, it exhibits significant overestimation in the Walker2d environment. These results collectively demonstrate the critical role of objective alignment in consistently ensuring conservative value estimation and preventing overestimation errors.
Table 2: Results on the D4RL Gym benchmark. The numbers reported for DAMO are the normalized scores averaged over the final iteration of training across 4 seeds, with $\pm$ standard deviation. Since we haven’t obtained the results of O-DICE on random datasets, we didn’t include it in the table and used a dash (-) to indicate the absence of data. The top three scores are highlighted in bold. We use an asterisk (*) for the best score, an underscore (_) for the second, and bold text for the third.
Task Name Algae Opti O-DICE MOPO COMBO TT RAMBO MOBILE DAMO (Ours)
halfcheetah-random -0.3 11.6 - 38.5 38.8 6.1 $\textbf{39.5}^{*}$ 38.3 $34.0\pm 3.0$
hopper-random 0.9 11.2 - $\textbf{31.7}^{*}$ 17.9 6.9 25.4 25.5 31.5 $\pm$ 0.4
walker2d-random 0.5 9.9 - 7.4 7.0 5.9 0.0 $\textbf{21.6}^{*}$ 0.0
halfcheetah-medium -2.2 38.2 47.4 72.4 54.2 46.9 $\textbf{77.9}^{*}$ 69.3 71.3 $\pm$ 0.5
hopper-medium 1.2 94.1 86.1 62.8 97.2 67.1 87.0 $\textbf{102.7}^{*}$ $93.8\pm 8.2$
walker2d-medium 0.3 21.8 84.9 84.1 81.9 81.3 84.9 $\textbf{88.7}^{*}$ $83.1\pm 8.0$
halfcheetah-medium-replay -2.1 39.8 44.0 $\textbf{72.1}^{*}$ 55.1 44.1 68.7 71.4 61.6 $\pm$ 3.0
hopper-medium-replay 1.1 36.4 $\textbf{99.9}^{*}$ 92.7 89.5 99.4 99.5 57.7 $\textbf{99.9}\pm\textbf{0.9}^{*}$
walker2d-medium-replay 0.6 21.6 83.6 85.9 56.0 82.6 89.2 80.5 89.4 $\pm$ $\textbf{3.0}^{*}$
halfcheetah-medium-expert -0.8 91.1 93.2 83.6 90.0 95.0 95.4 101.4 103.1 $\pm$ 0.9
hopper-medium-expert 1.1 111.5 110.8 74.6 111.0 110.0 88.2 107.7 112.3 $\pm$ $\textbf{3.6}^{*}$
walker2d-medium-expert 0.4 74.8 110.8 108.2 103.3 101.9 56.7 $\textbf{113.0}^{*}$ 109.4 $\pm$ 0.8
average score 0.0 43.4 - 68.7 66.0 61.1 66.1 72.2 $\textbf{74.4}^{*}$
6.2 Comparison Results
Dataset. We evaluate DAMO’s performance using the D4RL benchmark (Fu et al., 2020) within the MuJoCo simulation environment. Our comprehensive assessment covers 12 distinct datasets across three environments (HalfCheetah, Hopper, and Walker2d) and four dataset categories: random, medium, medium-replay, and medium-expert. Following standard evaluation protocols, we exclusively utilize the ”v2” versions of all datasets for consistent comparison.
Baseline. We compare DAMO with two types of offline RL baselines: DICE-based algorithms which follow a similar form of training objective with DAMO and Model-based algorithms which achieve SOTA performance. For DICE-based approaches, Algae-DICE (Nachum et al., 2019b) transforms the intractable state-action-level constraint into a unified objective for policy training. Opti-DICE (Lee et al., 2021) directly estimates the stationary distribution corrections of the optimal policy to attain a high-rewarding policy. O-DICE (Mao et al., 2024a) using the orthogonal-gradient update to diminish the confliction of different gradients during policy training. For model-based approaches, MOPO (Yu et al., 2020) penalizes rewards via the predicted variance of ensembled dynamics models. COMBO (Yu et al., 2021) implements CQL within a model-based framework. TT (Janner et al., 2019) uses a transformer to model offline trajectories and employs beam search for planning. RAMBO (Rigter et al., 2022) adversarially trains the policy and the dynamics model within a robust framework. MOBILE (Sun et al., 2023) penalizes the value learning of synthetic data based on the estimated uncertainty of the Bellman Q-function, derived from an ensemble of models.
Comparison Results. Table 2 reports the scores in D4RL benchmark. Overall, DAMO demonstrates superior performance, achieving the highest average score across all baseline methods. Notably, for high-quality datasets such as medium-expert and medium-replay, DAMO consistently attains state-of-the-art (SOTA) or near-SOTA performance. These results suggest that effective OOD constraints, combined with a high-quality behavior policy, enable the agent to learn high-reward policies within in-distribution (ID) regions. Furthermore, DAMO maintains strong performance on medium-quality datasets. This demonstrates the robustness of DAMO in handling less-than-ideal data conditions and highlights the importance of ensuring both model-environment policy consistency and synthetic-offline data compatibility. However, the performance of DAMO on random datasets reveals limitations, indicating the need for more aggressive exploration strategies in low-quality data scenarios to identify regions beyond those covered by the offline dataset. This phenomenon is particularly pronounced in the Walker2d-random dataset, where DAMO struggles to learn an effective policy. One possible explanation is that the optimization landscape of the environment may contain numerous suboptimal solutions corresponding to saddle points in the maximin framework. If this is the case, such structural characteristics could potentially make DAMO more susceptible to becoming trapped in local optima, thereby limiting its ability to discover better solutions. However, further investigation is needed to confirm this hypothesis.
6.3 Effects of Coefficient $\alpha$
(a) Tuning of $\alpha$
(b) DAMO without fixing $\alpha$
Figure 5: Two sets of experiments on hyperparameter $\alpha$. (a) The tuning experiment of $\alpha$ was performed on the hopper-medium-expert dataset. (b) The fixing experiment of $\alpha$ was performed on the halfcheetah-medium-expert dataset for validation.
DAMO is regulated by a key hyperparameter: the hyperparameter $\alpha$ which controls the extent of policy discrepancies. In practical implementation, we fix $\alpha$ in the actor training. While we consider $\alpha$ as a hyperparameter controlling the degree of conservatism, the empirical results in Fig. 5(a) demonstrate that DAMO maintains robust performance across a wide range of $\alpha$ values, with its primary impact observed in the final stage performance. Notably, DAMO achieves superior performance with larger $\alpha$ values (e.g., $\alpha=3.0,5.0$), whereas insufficiently small values (e.g., $\alpha=1.0$) tend to promote excessive risk-taking during exploration, ultimately failing to converge to an optimal policy. To enhance training stability, we initially fix $\alpha$ to 1.0 in the actor training. However, experimental results demonstrate that DAMO achieves competitive performance even without this specific implementation. To further validate this observation, we conduct comparative experiments on the HalfCheetah environment, evaluating both the standard DAMO implementation and its variant without fixed $\alpha$ under their respective optimal hyperparameter configurations. As shown in Fig. 5(b), both settings achieve a competitive performance, confirming that DAMO maintains robust performance regardless of the implementation of fixing $\alpha$.
7 Conclusion and Discussion
This paper critically examines limitations in existing offline model-based RL approaches, highlighting the crucial need to address behavior inconsistency caused by policy discrepancies between biased models and real environments. Building upon our analysis, we introduce a unified framework that concurrently ensures model-environment policy consistency and synthetic-offline data compatibility, with comprehensive experimental validation demonstrating its effectiveness. Significantly, the maximin optimization framework of DAMO demonstrates the benefits of aligning the actor and critic training objectives within the actor-critic architecture, ensuring essential conservatism during policy updates. The limitations of our work are focused on two-fold: the sampling of the initial state and the computing of the data alignment term. Although we assume the distribution of the initial state is known, the offline dataset does not explicitly label which states are initial states. Therefore, we randomly sample a state from the offline dataset to serve as the initial state, which creates a gap between theory and practice. When calculating the data alignment term, we trained a classifier to assist in the process. However, the classifier’s accuracy is critical: imprecise classifiers fail to estimate the modified reward accurately, leading to unstable training and potentially invalidating the practical objective as a lower bound. Future work could extend DAMO to areas such as offline-to-online RL (Li et al., 2023), where methods with strong adaptability and the ability to bridge the synthetic-to-real distribution gap are in high demand.
Impact Statement
This work presents an effective solution for addressing the distribution shift in offline model-based RL through dual policy-environment alignment, enabling robust simulation-to-reality transfer in safety-critical domains such as medical robotics. While advancing certifiable policy deployment via implicit behavior constraints and adaptive value estimation, challenges remain in reward specification sensitivity and residual model bias mitigation. The framework highlights the critical need for continuous dynamic alignment calibration to ensure reliable real-world RL system performance.
Appendix A Proofs in the Main Text
We present a detailed proof process for each theorem discussed in the paper. The sequence of proofs follows the logic of derivation rather than strictly adhering to their order of appearance in the main text.
Theorem A.1.
The KL-divergence has the following upper bound:
$D_{KL}(\rho_{T}^{\pi}\|\rho_{T}^{\pi_{\beta}})\leq\mathrm{E}_{\rho_{T}^{\pi}}[\log\dfrac{\rho_{T}^{\pi}}{\rho_{T}^{\pi_{\beta}}}]+D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi}).$
Proof.
$D_{KL}(\rho_{T}^{\pi}\|\rho_{T}^{\pi_{\beta}})$ $=\mathrm{E}_{\rho_{T}^{\pi}}[\log\dfrac{\rho_{T}^{\pi}}{\rho_{T}^{\pi_{\beta}}}]$
$=\mathrm{E}_{\rho_{T}^{\pi}}[\log\dfrac{\rho_{T}^{\pi}}{\rho_{T}^{\pi_{\beta}}}\cdot\dfrac{\rho_{M}^{\pi}}{\rho_{M}^{\pi}}]$
$=\mathrm{E}_{\rho_{T}^{\pi}}[\log\dfrac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}]+D_{KL}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})$
$\leq\mathrm{E}_{\rho_{T}^{\pi}}[\log\dfrac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}]+D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})$
Corollary A.2.
We have the following surrogate objective
$\arg\max\limits_{\pi}\mathrm{E}_{\rho_{T}^{\pi}}[r(s,a,s^{\prime})-\log\dfrac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}]-D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})$
A discounted factor $\alpha$ will be added for training stability.
$\arg\max\limits_{\pi}\mathrm{E}_{\rho_{T}^{\pi}}[r(s,a,s^{\prime})-\alpha\log\dfrac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}]-\alpha D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})$ (5)
Lemma A.3.
We can represent f-divergence in a function form: $D_{f}(x\|p)=\mathbb{E}_{z\sim p}f(\frac{x}{p})$, then its Fenchel conjugate function is $\mathbb{E}_{z\sim p}f_{\star}(y(z))$
Proof.
We use $g_{\star}(x)$ to denote $D_{f}(x\|p)$
$g_{\star}(x)$ $=\sup\limits_{y\in\mathcal{X}}\{\langle x,y\rangle-D_{f}(y\|p)\}$
[…]
$=\int_{R}p(z)\sup\limits_{z}[x(z)\frac{y(z)}{p(z)}-f(\frac{y(z)}{p(z)})]dz$
$=\int_{R}p(z)\sup\limits_{z}[x(z)y(z)-f(y(z))]dz$
[…]
$=\mathrm{E}_{z\sim p}f_{\star}(y(z))$
Corollary A.4.
For f-divergence $D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})$
$D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})=\max\limits_{y(s,a,s^{\prime})}\mathrm{E}_{\rho_{T}^{\pi}}[y(s,a,s^{\prime})]-\mathrm{E}_{\rho_{M}^{\pi}}[f_{\star}(y(s,a,s^{\prime}))]$
Proof.
Verifying $D_{f}$ is a convex, closed function. A convex and closed function’s double conjugate function is itself, so $D_{f}$ can be represented in a MAX form. ∎
Theorem A.5.
Surrogate objective (5) is equivalent to the following maximin problem
$\max\limits_{\pi}\min\limits_{y(s,a,s^{\prime})}\mathrm{E}_{\rho_{T}^{\pi}}[r(s,a,s^{\prime})-\alpha\log\dfrac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}-\alpha y(s,a,s^{\prime})]+\alpha\mathrm{E}_{\rho_{M}^{\pi}}[f_{\star}(y(s,a,s^{\prime}))]$
Proof.
According to lemma A.4, the objective can be presented in a Min form. ∎
Theorem A.6.
The following objective is equivalent to the surrogate objective
$\max\limits_{\pi}\min\limits_{Q(s,a)}J(\pi,Q(s,a))=\max\limits_{\pi}\min\limits_{Q(s,a)}(1-\gamma)\mathrm{E}_{s\sim\mu_{0},a\sim\pi}[Q(s,a)]+\mathrm{E}_{\rho_{M}^{\pi}}f_{\star}(\Phi(s,a,s^{\prime})/\alpha)$ (6)
Proof.
Defining $y(s,a,s^{\prime})=\dfrac{r(s,a,s^{\prime})-\alpha\log\frac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}+\tau^{\pi}Q(s,a)-Q(s,a)}{\alpha}$. For the convenience, we define $\beta_{t}(s,a,s^{\prime})$
$\beta_{t}(s,a,s^{\prime})=Pr[s_{t}=s,a_{t}=a,s_{t+1}=s^{\prime}|s_{0}\sim\mu_{0},a_{t}\sim\pi,s_{t+1}\sim T]$
Then, $\rho_{T}^{\pi}(s,a,s^{\prime})=(1-\gamma)\sum\limits_{t=0}^{\infty}\gamma^{t}\beta_{t}$. Similarly, we can define $\beta_{t}(s,a,s^{\prime},a^{\prime})$.
Substituting $y(s,a,s^{\prime})$ with $[r(s,a,s^{\prime})-\alpha\log\frac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}+\tau^{\pi}Q(s,a)-Q(s,a)]/\alpha$ and we have:
$\max\limits_{\pi}\min\limits_{Q(s,a)}$ $\mathrm{E}_{\rho_{T}^{\pi}}[Q(s,a)-\tau^{\pi}Q(s,a)]$
$=$ $\sum\limits_{s,a,s^{\prime}}\rho_{T}^{\pi}(s,a,s^{\prime})[Q(s,a)-\gamma\sum\limits_{a^{\prime}}Q(s^{\prime},a^{\prime})\pi(a^{\prime}|s^{\prime})]$
$=$ $\sum\limits_{s,a,s^{\prime}}[\rho_{T}^{\pi}(s,a,s^{\prime})Q(s,a)]-\sum\limits_{s,a,s^{\prime},a^{\prime}}[\gamma\rho_{T}^{\pi}(s,a,s^{\prime})Q(s^{\prime},a^{\prime})\pi(a^{\prime}|s^{\prime})]$
$=$ $(1-\gamma)[\sum\limits_{s,a,s^{\prime}}\sum\limits_{t=0}^{\infty}\gamma^{t}\beta_{t}(s,a,s^{\prime})Q(s,a)-\sum\limits_{s,a,s^{\prime},a^{\prime}}\sum\limits_{t=0}^{\infty}\gamma^{t+1}\beta_{t}(s,a,s^{\prime})Q(s^{\prime},a^{\prime})\pi(a^{\prime}|s^{\prime})]$
$=$ $(1-\gamma)[\sum\limits_{s,a,s^{\prime}}\sum\limits_{t=0}^{\infty}\gamma^{t}\beta_{t}(s,a,s^{\prime})Q(s,a)-\sum\limits_{s,a,s^{\prime},a^{\prime}}\sum\limits_{t=0}^{\infty}\gamma^{t+1}\beta_{t}(s,a,s^{\prime},a^{\prime})Q(s^{\prime},a^{\prime})]$
$=$ $(1-\gamma)[\sum\limits_{s,a}\sum\limits_{t=0}^{\infty}\gamma^{t}\beta_{t}(s,a)Q(s,a)-\sum\limits_{s^{\prime},a^{\prime}}\sum\limits_{t=0}^{\infty}\gamma^{t+1}\beta_{t}(s^{\prime},a^{\prime})Q(s^{\prime},a^{\prime})]$
$=$ $(1-\gamma)\sum\limits_{s,a}\beta_{0}(s,a)Q(s,a)$
$=$ $(1-\gamma)\mathrm{E}_{s\sim\mu_{0},a\sim\pi}[Q(s,a)]$
Theorem A.7.
The optimal solution of maximin question (6)’s inner question $Q^{\star}(s,a)$ satisfies
$Q^{\star}(s,a)=r(s,a,s^{\prime})-\alpha\log\dfrac{\rho^{\pi}_{M}}{\rho^{\pi_{\beta}}_{T}}-\alpha f^{\prime}(\dfrac{\rho^{\pi}_{T}}{\rho^{\pi}_{M}})+\tau^{\pi}Q^{\star}(s,a)$ (7)
Proof.
$f(x)$ is a convex function, so its conjugate function $f_{\star}$ is also convex. The objective function below is convex about $y(s,a,s^{\prime})$
$\max\limits_{\pi}\min\limits_{y(s,a,s^{\prime})}\mathbb{E}_{\rho_{T}^{\pi}}[r(s,a,s^{\prime})-\alpha\log\dfrac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}-\alpha y(s,a,s^{\prime})]+\alpha\mathbb{E}_{\rho_{M}^{\pi}}[f_{\star}(y(s,a,s^{\prime}))]$
Then the optimal solution of inner question is $y^{\star}(s,a,s^{\prime})=(f^{\prime}_{\star})^{-1}(\dfrac{\rho^{\pi}_{T}}{\rho^{\pi}_{M}})$. For any convex function, we have $(f^{\prime}_{\star})^{-1}(x)=f^{\prime}(x)$, so $y^{\star}(s,a,s^{\prime})=f^{\prime}(\dfrac{\rho^{\pi}_{T}}{\rho^{\pi}_{M}})$.
Recalling that we employ the variable substitution:$y(s,a,s^{\prime})=[r(s,a,s^{\prime})-\alpha\log\frac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}+\tau^{\pi}Q(s,a)-Q(s,a)]/\alpha$, so we have
$y^{\star}(s,a,s^{\prime})=\dfrac{r(s,a,s^{\prime})-\alpha\log\frac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}+\tau^{\pi}Q^{\star}(s,a)-Q^{\star}(s,a)}{\alpha}$
Then the optimal solution $Q^{\star}(s,a)$ satisfies
$Q^{\star}(s,a)=r(s,a,s^{\prime})-\alpha\log\dfrac{\rho^{\pi}_{M}}{\rho^{\pi_{\beta}}_{T}}-\alpha f^{\prime}(\dfrac{\rho^{\pi}_{T}}{\rho^{\pi}_{M}})+\tau^{\pi}Q^{\star}(s,a)$
Appendix B OOD Issues under Dynamics Model Integration
OOD issues in offline model-based RL are intrinsically linked to the mismatch between synthetic and real data distributions, compounded in model-based settings, where model inaccuracies introduce additional sources of error.
B.1 OOD Actions
The dynamics model generates data that can significantly deviate from the distribution of the offline dataset, leading to OOD actions. This triggers negative model exploitation (Levine et al., 2020), resulting in poor policy performance. Aligning the distributions of synthetic and offline data can effectively mitigate these issues. The model, $M$, is used to generate synthetic data for training, but it is essential to avoid taking OOD actions when rolling out policy in the model, to ensure accurate value estimation.
Specifically, in an offline setting, where interaction with the real environment is not possible, overestimating the value of OOD actions leads to model exploitation. Consequently, the agent may learn policies that perform poorly in the underlying environment compared to the model, leading to algorithmic failure during deployment. To avoid model exploitation, it is crucial to align synthetic with offline data. Given offline data collected under the behavior policy $\pi_{\beta}$, We aim to find a policy $\pi$ that behaves in model $M$ like $\pi_{\beta}$ does in environment $T$. Specifically, we aim to match the distribution of $(s,a)_{M}^{\pi}$ with $(s,a)_{T}^{\pi_{\beta}}$, ensuring accurate value estimation of $(s,a)_{M}^{\pi}$.
B.2 OOD states
The performance of the learning policy during deployment is critically dependent on the alignment of transition pairs between the learned model and the real environment. Significant discrepancies between these transitions can cause performance degradation and lead to OOD states, highlighting the necessity of correcting transition biases during training. Even if the learned policy performs well in the model, it is important, during real-environment deployment, to avoid OOD states, for which we cannot accurately estimate the value function in the offline setting.
For example, if the learned model predicts the transition $(s,a,s_{1})$, but the real environment produces a transition $(s,a,s_{2})$, where $s_{1}$ is a high-value state and $s_{2}$ is a low-value state, the policy will perform poorly in the real environment. The situation worsens if $s_{2}$ is an OOD state that has never been encountered in model $M$ before. In this scenario, the agent cannot estimate the value of the state-action pair $(s_{2},a^{\prime})$, leading to ineffective decision-making and model-environment policy inconsistency. In an offline setting, since we cannot adjust the policy by interacting with the environment, we must approximate and correct the discrepancies between the model $M$ and the real environment $T$ during training. This ensures that the distribution of transition $(s,a,s^{\prime})_{M}^{\pi}$ matches the distribution of $(s,a,s^{\prime})_{T}^{\pi}$ in the underlying environment.
In summary, while optimizing the policy $\pi$ to maximize returns, we should constrain it in two directions: 1) aligning the synthetic data $(s,a,s^{\prime})_{M}^{\pi}$ with the offline data $(s,a,s^{\prime})_{T}^{\pi_{\beta}}$, and 2) ensuring that $\pi$ exhibits consistent behavior in both dynamics models and real environments.
Appendix C Implementation Details
In this section, we present the detailed implementation of DAMO.
C.1 Selection of f-divergence
As established in Theorem 5.3, the f-divergence must consistently exceed the KL-divergence to guarantee that the objective in Equation (2) serves as a lower bound for the standard RL objective $\mathcal{J}(\pi)$. Based on this, we specifically choose $f(x)=\frac{1}{3}(x-1)^{3}$, and its conjugate function is $f_{\star}(x)=\frac{2}{3}(x-1)^{\frac{3}{2}}.$ In principle, any choice of f-divergence is permissible, and we adhere to the selection made in the prior work (Luo et al., 2024b). Due to the tight schedule, we haven’t tried other choices of f-divergence. However, we suppose that the choice of f-divergence won’t enormously affect the effectiveness of DAMO, since f-divergence primarily serves as a measure of discrepancy.
Table 3: Hyperparameter of DAMO when optimizing policy
Hyperparameter Value Description
$n$ 7 Ensemble number of dynamics model
$n_{elite}$ 5 Number of dynamics model for prediction
$d$ 0.5 Sampling ratio for classifier training
$\gamma$ 0.99 Discounted factor.
$l_{r}$ of critic $3\times 10^{-4}$ Critic learning rate.
Batch size 256 Batch size for each update.
$N_{iter}$ 3M Total gradient steps.
C.2 Modification of actor training objective
When optimizing the actor network using objective (2), we observed that the substantial residual of the Bellman error often disrupts the optimization of the actor and leads to training instability. To stabilize the training process, we followed the implementation of previous DICE work (Mao et al., 2024a; Sikchi et al., 2023) and fixed the hyperparameter $\alpha$ to $1.0$ during the optimization of the actor network. The specific form of the actor training objective is as follows:
$\max\limits_{\pi}(1-\gamma)\mathrm{E}_{s\sim\mu_{0},a\sim\pi}[\tilde{Q}(s,a)]\\ +\mathrm{E}_{\rho_{M}^{\pi}}[f_{\star}(\Phi(s,a,s^{\prime}))]$
It should be noted that the effectiveness of DAMO does not rely on this trick. We demonstrated this in the experimental section.6.3.
C.3 Model, classifier and Policy Optimization
For dynamics model $M$, it was depicted as a probabilistic neural network that generates a Gaussian distribution for the subsequent state and reward, contingent upon the current state and action:
$m_{\theta}(s_{t+1},r_{t}|s_{t},a_{t})=\mathcal{N}(\mu_{\theta}(s_{t},a_{t}),\Sigma_{\theta}(s_{t},a_{t})).$
Our model training approach is consistent with the methodology used in prior works (Yu et al., 2020; Sun et al., 2023). We train an ensemble of seven dynamics models and select the best five based on their validation prediction error from a held-out set containing 1000 transitions in the offline dataset $D_{R}$. Each model in the ensemble is a 4-layer feedforward neural network with 200 hidden units. We randomly choose one model from the best five models during model rollouts.
For classifier $h_{\psi}$, we implement the following approach: At each gradient step, we randomly sample data from both offline dataset $D_{R}$ and synthetic dataset $D_{M}$ with a fixed ratio $d$. This method ensures a balanced utilization of data from both DL and DG during training.
Policy optimization is conducted using the Actor-Critic framework in our approach, with hyperparameter configurations aligned with previous offline model-based settings. During each update, a batch of 256 transitions is sampled, where 5% is derived from the offline dataset $D_{R}$ and the remaining 95% from the synthetic dataset $D_{M}$. Additionally, We use two common tricks in DAMO: double Q-learning and entropy regularizer. The specific hyperparameter settings applied for the D4RL benchmark are detailed in Table.3
Appendix D Experiment Setting
D.1 Benchamark
We evaluate our approach using the D4RL benchmark (Fu et al., 2020), focusing on the “v2” version of Gym tasks. For DICE-based methods (Algae-DICE, Opti-DICE, and O-DICE), we directly adopt the scores reported in their respective original papers. However, since the results for O-DICE on the random dataset are not provided in the original paper, we exclude them from our comparison. For offline model-based methods, we retrained MOPO (Yu et al., 2020) and MOBILE (Sun et al., 2023) on the “v2” Gym datasets. For the remaining methods, we directly report the scores provided in their original papers. All results are summarized in Table 2.
Table 4: Hyperparameters of DAMO used in the D4RL datasets.
Task Name $\alpha$ $K$ $l_{r}$ of actor
halfcheetah-random 0.1 5 $1\times 10^{-4}$
hopper-random 2 5 $1\times 10^{-5}$
walker2d-random 2 5 $1\times 10^{-4}$
halfcheetah-medium 0.5 5 $1\times 10^{-4}$
hopper-medium 2.0 5 $1\times 10^{-5}$
walker2d-medium 2 1 $5\times 10^{-5}$
halfcheetah-medium-replay 0.1 5 $1\times 10^{-4}$
hopper-medium-replay 3 5 $1\times 10^{-4}$
walker2d-medium-replay 1.5 5 $1\times 10^{-4}$
halfcheetah-medium-expert 2.5 5 $1\times 10^{-4}$
hopper-medium-expert 3 5 $1\times 10^{-5}$
walker2d-medium-expert 1.5 3 $1\times 10^{-4}$
D.2 Hyperparameters
We provide a list of the hyperparameters that were tuned during our experiments. The detailed configurations are summarized in Table 4.
Coefficient $\alpha$. The coefficient $\alpha$ acts as the sole hyperparameter in our objective function, regulating the divergence between the training policy and the behavior policy. We tune $\alpha$ in the range of $[0.1,3.0].$
Rollout Length $K$. Similar to MOPO, we employ short-horizon branch rollouts in DAMO. For Gym tasks in the D4RL benchmark, the horizon parameter $K$ is tuned within the range of $\{1,3,5\}$.
The actor learning rate $l_{r}$ Since DAMO operates as a Maximin optimization problem and is susceptible to saddle points, we optimize the actor’s learning rate to achieve robust performance. The choice of learning rate plays a pivotal role in DAMO’s effectiveness. We tune the actor learning rate $l_{r}$ in the range of $\{1\times 10^{-5},5\times 10^{-5},1\times 10^{-4}\}$.
D.3 Computing Infrastructure
The experimental setup includes a GeForce GTX 3090 GPU and an Intel(R) Xeon(R) Gold 6330 CPU at 2.00GHz. The implementation is built upon the OfflineRL-Kit library, with DAMO adopting the software libraries and frameworks specified therein.

### v3

@@
-Dual Alignment Maximin Optimization for Offline Model-based RL
-Abstract
+Dual Alignment Maximin Optimization for Offline MBRL
+Abstract.
 Offline reinforcement learning agents face significant deployment challenges due to the synthetic-to-real distribution mismatch.
-While most prior research has focused on improving the fidelity of synthetic sampling and incorporating off-policy mechanisms, the directly integrated paradigm often fails to ensure consistent policy behavior in biased models and underlying environmental dynamics, which inherently arise from discrepancies between behavior and learning policies.
-In this paper, we first shift the focus from model reliability to policy discrepancies while optimizing for expected returns, and then self-consistently incorporate synthetic data, deriving a novel actor-critic paradigm, Dual Alignment Maximin Optimization (DAMO).
-It is a unified framework to ensure both model-environment policy consistency and synthetic and offline data compatibility.
+While most prior research has focused on improving the fidelity of synthetic sampling and incorporating off-policy mechanisms, the directly integrated paradigm often fails to ensure consistent policy behavior in biased models and underlying environmental dynamics, which arise from discrepancies between behavior and learning policies.
+In this paper, we first shift the focus from model reliability to constraining policy discrepancies while optimizing for expected returns, and then self-consistently incorporate synthetic data, deriving a novel maximin paradigm, Dual Alignment Maximin Optimization (DAMO).
+DAMO is a unified framework to ensure both model-environment policy consistency and synthetic and offline data compatibility.
 The inner minimization performs dual conservative value estimation, aligning policies and trajectories to avoid out-of-distribution states and actions, while the outer maximization ensures that policy improvements remain consistent with inner value estimates.
-Empirical evaluations demonstrate that DAMO effectively ensures model and policy alignments, achieving competitive performance across diverse benchmark tasks.
-Machine Learning, Offline Reinforcement learning
-1 Introduction
-Offline reinforcement learning (RL) (Lange et al., 2012; Levine et al., 2020) aims to learn policies from a pre-collected dataset generated by a behavioral policy within the real environment.
-This paradigm helps avoid the safety risks and high costs associated with direct interactions, making RL applicable in real-life scenarios, such as healthcare decision-making support and autonomous driving (Emerson et al., 2023; Sinha et al., 2022; Mnih et al., 2015).
-Offline model-based RL (Yu et al., 2020; Kidambi et al., 2020; Sun et al., 2023) improves upon this by training a dynamics model and using it to generate synthetic data for policy training or planning.
-The introduction of dynamics models enhances the sample efficiency and allows the agent to answer counterfactual queries (Levine et al., 2020).
-However, despite strong performance in learned models, policies often degrade significantly when deployed in real environments due to mismatches between synthetic and real distributions.
-Most prior works have focused on enhancing the reliability of synthetic samplings from learned models, and directly incorporating off-policy optimization methods, such as SAC (Haarnoja et al., 2018), to address this mismatch.
-The representative approach is MOPO (Yu et al., 2020), which quantifies model uncertainty by measuring the predicted variance of learned models and subsequently penalizes model-generated trajectories with high uncertainty.
-This methodology has inspired subsequent research like MOReL (Kidambi et al., 2020) and MOBILE (Sun et al., 2023), which adopt similar uncertainty-aware frameworks.
-These approaches help generate synthetic data that remains compatible with the offline dataset and construct conservative value functions to mitigate out-of-distribution (OOD) actions.
-However, creating a perfect learned model from limited offline datasets is impossible, and as illustrated in Fig.
-1, simply combining these methods with off-policy mechanisms still fails to resolve the inherent discrepancies in policy behaviors between the learned model and the underlying environment, leading to inconsistent policy performance and leaving the distribution shift problem unsolved.
-In this paper, we trace the inherent root of distribution shift challenges in offline RL to the discrepancies between behavior and learning policies in the underlying environmental dynamics.
-While optimizing for expected returns, we introduce a regularized policy optimization objective for offline RL, that constrains the visitation distribution discrepancies between the behavioral and the learning policies in real environments.
-We further self-consistently incorporate synthetic data into this objective, deriving a novel maximin optimization objective for offline model-based RL.
-Building on these objectives, we propose Dual Alignment Maximin Optimization (DAMO), a consistent actor-critic framework that ensures both coherent alignments between model and environment behavior and model-generated and offline data.
-DAMO iteratively applies maximin optimization in two phases.
-In the inner minimization phase, equivalent to critic training, the focus is on the dual conservative value estimation.
-This value estimation implicitly harmonizes learning policy behaviors in both the model and the environment to avoid OOD states, while explicitly aligning synthetic and offline data to prevent OOD actions.
-In the outer maximization phase, corresponding to actor learning, policy improvements are guided by the inner value estimates through objective alignment.
-This preserves reliably conservative behavior and effectively mitigates OOD issues throughout both policy optimization and deployment.
-Additionally, DAMO incorporates classifiers to approximate the data alignment terms, facilitating practical implementation.
-Extensive experiments showcase that DAMO successfully mitigates OOD issues and achieves competitive performance across various benchmarks.
-Our contributions are summarized as the following:
-  • We derive a consistent maximin object for offline model-based RL, ensuring both alignments between model and environment policy behaviors, and synthetic and offline data.
-  • We develop DAMO, a novel actor-critic framework that employs maximin optimization to achieve the dual conservative value estimation and the consistent policy improvement.
-  • We conduct extensive experiments to validate that DAMO effectively alleviates the distribution shift challenge, delivering superior performance across varying benchmark tasks.
+Empirical evaluations demonstrate that DAMO effectively ensures model and policy alignments, achieving competitive performance across diverse offline benchmark tasks.
+Key words and phrases:
+Offline Model-Based RL, OOD Issues, Maximin Optimization
+1. Introduction
+Offline reinforcement learning (RL) Lange et al.
+(2012); Levine et al.
+(2020) aims to learn policies from a pre-collected dataset generated by a behavior policy within the real environment.
+This paradigm helps avoid the potential safety risks and high costs associated with direct online interactions, making RL applicable in real-life scenarios, such as healthcare decision-making support and autonomous driving Emerson et al.
+(2023); Sinha et al.
+(2022); Mnih et al.
+(2015).
+Offline model-based RL (MBRL) Yu et al.
+(2020); Kidambi et al.
+(2020); Sun et al.
+(2023) improves upon this by training a dynamics model and using it to generate synthetic data for policy training.
+The introduction of the dynamics model enhances the sample efficiency and allows the agent to answer counterfactual queries Levine et al.
+(2020).
+However, despite strong performance in dynamics models, trained policies often degrade when deployed in real scenarios due to mismatches between synthetic and real distributions.
+Most prior works have focused on enhancing the reliability of synthetic samplings from dynamics models, and directly incorporating off-policy optimization methods such as SAC Haarnoja et al.
+(2018), to address this mismatch.
+The representative approach is MOPO Yu et al.
+(2020), which quantifies dynamics model uncertainty by measuring the predicted variance of learned models and penalizes synthetic trajectories with high uncertainty.
+This methodology has inspired subsequent researchs like MOReL Kidambi et al.
+(2020) and MOBILE Sun et al.
+(2023), which adopt similar uncertainty-aware frameworks.
+These approaches construct conservative value functions to mitigate overestimation of out-of-distribution (OOD) actions, thereby reducing execution of these unreliable actions and facilitating the generation of synthetic data compatible with the offline dataset.
+However, as illustrated in Fig.
+1, although this uncertainty-aware framework aligns the synthetic data with offline data and mitigates OOD actions during policy training, it fails to alleviate OOD trajectories when deployed into real scenarios, leading to inconsistent policy behavior during deployment and leaving the OOD issues in offline MBRL unresolved.
+To bridge this gap, we start from a regularized policy optimization objective, which limits the discrepancy between the visitation distribution of the training policy and the behavior policy under the real environment.
+By self-consistently incorporating the synthetic data into the regularized objective, we derive a unified maximin optimization objective for offline MBRL.
+Building on this objective, we propose Dual Alignment Maximin Optimization (DAMO), a unified actor-critic framework to ensure dual conservative estimation.
+DAMO iteratively applies maximin optimization in two phases: the inner minimization and the outer maximization.
+In the inner minimization, equivalent to critic training, we provide a theoretical analysis showing that it enforces dual conservative estimation.
+This value estimation explicitly aligns the synthetic and offline data to prevent OOD actions during training, while implicitly aligning learning policy behaviors in both the learned model and the real environment to avoid OOD states during deployment.
+In the outer maximization, corresponding to actor learning, policy improvements are guided by the inner value estimates through objective alignment.
+This preserves conservative behavior and mitigates OOD issues throughout both policy optimization and deployment.
+Additionally, DAMO incorporates a classifier to enable the dual conservative estimation during practical implementation.
+Extensive experiments showcase that DAMO successfully mitigates OOD issues and achieves competitive performance across various offline benchmarks.
+Our contributions are summarized as follows:
+  • We derive a consistent maximin objective for offline MBRL, ensuring both alignments between dynamics model and real environment policy behaviors, and synthetic and offline data.
+  • We develop DAMO, an actor-critic framework that employs maximin policy optimization to achieve both the dual conservative value estimation and the consistent policy improvement.
+  • We conduct experiments to validate that DAMO alleviates the OOD issues throughout training and deployment, delivering superior performance across varying benchmarks.
 (a) MOPO
@@
 (c) DAMO
-Figure 1: The distribution of transition pair $(s,a,s^{\prime})$ across three datasets (Offline Data, Real Data, and Synthetic Data) are shown for the hopper-medium-expert task, comparing two SOTA methods (MOPO and MOBILE) with our proposed DAMO framework.
-While MOPO and MOBILE align synthetic and offline data well (indicated by a high overlap), they fail to address OOD states in the real environment (circled by an ellipse drawn with a green dotted line).
+Figure 1.
+The distributions of transition pair $(s,a,s^{\prime})$ across three datasets (Offline Data, Real Data, and Synthetic Data) are shown for the hopper-medium-expert task, comparing MOPO and MOBILE with our proposed DAMO framework.
+The Real Data is collected through executing the policy trained using specific offline methods in the real environment.
+While MOPO and MOBILE align synthetic and offline data well (indicated by a high overlap), they fail to mitigate OOD states in the real environment (circled by an ellipse).
 In contrast, DAMO effectively eliminates the OOD region.
-2 Related Work
-Offline Model-based RL trains a dynamics model using offline data to approximate environmental dynamics and performs conservative policy optimization (Lu et al., 2021) with model-generate data to mitigate OOD issues.
-MOPO (Yu et al., 2020) penalizes state-action pairs with high model uncertainty, effectively discouraging the agent from selecting OOD actions.
-MOReL (Kidambi et al., 2020) constructs a pessimistic Markov decision process to prevent the agent from entering OOD regions, ensuring safer and more reliable policy learning.
-COMBO (Yu et al., 2021) extends Conservative Q-learning (Kumar et al., 2020) to the model-based setting by regularizing the value function on OOD samples.
-RAMBO (Rigter et al., 2022) employs an adversarial training framework, optimizing both the policy and model jointly to ensure accurate transition predictions while maintaining robustness.
-MOBILE (Sun et al., 2023) incorporates penalties into value learning by utilizing uncertainty in Bellman function estimates derived from ensemble models.
-SAMBO (Luo et al., 2024a) imposes reward penalties and fosters exploration by incorporating model and policy shifts inferred through a probabilistic inference framework.
+2. Related Work
+Offline Model-based RL trains a dynamics model using offline data to approximate environmental dynamics and performs conservative policy optimization Lu et al.
+(2021) with model-generate data to mitigate OOD issues.
+MOPO Yu et al.
+(2020) penalizes state-action pairs with high model uncertainty, effectively discouraging the agent from selecting OOD actions.
+MOReL Kidambi et al.
+(2020) constructs a pessimistic Markov decision process to prevent the agent from entering OOD regions, ensuring safer and more reliable policy learning.
+COMBO Yu et al.
+(2021) extends Conservative Q-learning Kumar et al.
+(2020) to the model-based setting by regularizing the value function on OOD samples.
+RAMBO Rigter et al.
+(2022) employs an adversarial training framework, optimizing both the policy and model jointly to ensure accurate transition predictions while maintaining robustness.
+MOBILE Sun et al.
+(2023) incorporates penalties into value learning by utilizing uncertainty in Bellman function estimates derived from ensemble models.
+SAMBO Luo et al.
+(2024b) imposes reward penalties and fosters exploration by incorporating model and policy shifts inferred through a probabilistic inference framework.
 DICE-based Methods.
 Distribution Correction Estimation (DICE) is a technique that estimates stationary distribution ratios using duality theory.
-DICE has shown superior performance in evaluating discrepancies between distributions and has been widely applied in various RL domains, including off-policy evaluation (OPE) (Nachum et al., 2019a), offline imitation learning (IL) (Ma et al., 2022), and offline RL (Nachum et al., 2019b).
+DICE has shown superior performance in evaluating discrepancies between distributions and has been widely applied in various RL domains, including off-policy evaluation (OPE) Nachum et al.
+(2019a), offline imitation learning (IL) Ma et al.
+(2022), and offline RL Nachum et al.
+(2019b).
 In offline RL, DICE-based methods correct distribution shifts between offline data and environmental dynamics, formulating a tractable maximin optimization objective.
-For instance, AlgaeDICE (Nachum et al., 2019b) pioneers the application of DICE-based methods in offline RL, employing regularized dual objectives and Lagrangian techniques to address distribution shift challenges.
-OptiDICE (Lee et al., 2021) incorporates the Bellman flow constraint and directly estimates stationary distribution corrections for the optimal policy, eliminating the need for policy gradients.
-ODICE (Mao et al., 2024a) combines orthogonal-gradient updates with DICE to enforce state-action-level constraints.
-Unlike these model-free approaches, DAMO is, to the best of our knowledge, the first method to promote DICE insights in the model-based setting.
-While model-free methods typically handle discrepancies between two distributions, DAMO skillfully extends this concept by managing the differences across three distributions, using a divergence upper bound to ensure alignment.
-Figure 2: Intuiative examples for demystifying OOD issues in offline model-based RL.
+For instance, AlgaeDICE Nachum et al.
+(2019b) pioneers the application of DICE-based methods in offline RL, employing a regularized dual objective and Lagrangian techniques to address distribution shift challenges.
+OptiDICE Lee et al.
+(2021) incorporates the Bellman flow constraint and directly estimates stationary distribution corrections for the optimal policy, eliminating the need for policy gradients.
+ODICE Mao et al.
+(2024b) deploys orthogonalization techniques to eliminate the conflicts between varying gradients, and updates the policy with DICE to enforce state-action-level constraints.
+Unlike these proposed model-free approaches, DAMO is, to the best of our knowledge, the first method to promote DICE insights in the model-based setting.
+While model-free methods typically handle discrepancies between two distributions, DAMO skillfully extends this concept by managing the discrepancies across three distributions, using a divergence upper bound to ensure dual alignment.
+3. Preliminaries
+MDP.
+We consider the Markov decision process (MDP) defined by the six-element tuple $M=(\mathcal{S},\mathcal{A},T,r,\mu_{0},\gamma)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ denotes the action space, $T(s^{\prime}|s,a)$ is the environment transition, $r(s,a,s^{\prime})>0$ is the reward function, $\mu_{0}$ is the initial state distribution, and $\gamma\in(0,1)$ is the discounted factor.
+Given an MDP, the objective of RL is to find a policy $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})$ that maximizes the cumulative reward from the environment.
+This objective can be formally expressed as: $\pi^{\star}=\arg\max_{\pi}\mathbb{E}_{\mu_{0},\pi,T}[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t},s_{t+1})]$.
+In this paper, we utilize the dual form of the RL objective Puterman (2014), which is represented as follows:
+$\pi^{\star}=\arg\max\limits_{\pi}\mathbb{E}_{\rho^{\pi}_{T}}[r(s,a,s^{\prime})].$
+Here, $\rho^{\pi}_{T}(s,a,s^{\prime})$ is the transition occupancy measure, characterizing the probability distribution of transition pairs $(s,a,s^{\prime})$ induced by policy $\pi$ under the dynamics $T$ among the episode, defined by:
+$\rho_{T}^{\pi}(s,a,s^{\prime})=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}\mathbb{P}[s_{t}=s,a_{t}=a,s_{t+1}=s^{\prime}\\ |s_{0}\sim\mu_{0},a_{t}\sim\pi(\cdot|s_{t}),s_{t+1}\sim T(\cdot|s_{t},a_{t})].$
+$\mathbb{P}$ represents the probability. For simplicity, we will substitute $\rho^{\pi}_{T}(s,a,s^{\prime})$ with notation $\rho^{\pi}_{T}$ in the remainder of the article.
+4. Demystify OOD in Offline MBRL
+Figure 2.
+Intuitive examples for the OOD issue.
 The behavior policy $\pi_{\beta}$ primarily favors the top road ($a_{1}$) with a minor representation of the bottom road ($a_{3}$) and never selects the middle road ($a_{2}$).
-Consequently, the dynamics model lacks representation of medium road ($a_{2}$) and treats the $a_{3}$ as an OOD action.
+Thus, the dynamics model lacks representation of the $a_{2}$ and treats the $a_{3}$ as an OOD action.
 The policy $\pi$, trained on both offline and synthetic data, avoids $a_{3}$ when deployed in the real environment but struggles to evaluate $a_{2}$ due to its absence in $M$, resulting in occasionally getting into OOD state.
@@
 At the bottom, we show that while data alignment ensures a similar distribution between offline and synthetic data, it fails to guarantee behavior consistency between $M$ and $T$.
-3 Preliminaries
-MDP.
-We consider the Markov decision process (MDP) defined by the six-element tuple $M=(\mathcal{S},\mathcal{A},T,r,\mu_{0},\gamma)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ denotes the action space, $T(s^{\prime}|s,a)$ is the environment transition dynamics, $r(s,a,s^{\prime})>0$ is the reward function, $\mu_{0}$ is the initial state distribution, and $\gamma\in(0,1)$ is the discounted factor.
-Given an MDP, the objective of RL is to find a policy $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})$ that maximizes the cumulative reward from the environment.
-This objective can be formally expressed as: $\pi^{\star}=\arg\max_{\pi}\mathbb{E}_{\mu_{0},\pi,T}[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t},s_{t+1})]$.
-In the derivation of DAMO, we utilize the dual form of the RL objective (Puterman, 2014), which is represented as follows:
-$\pi^{\star}=\arg\max\limits_{\pi}\mathbb{E}_{\rho^{\pi}_{T}}[r(s,a,s^{\prime})].$
-Here, $\rho^{\pi}_{T}(s,a,s^{\prime})$ is the transition occupancy measure, characterizing the distribution of transition pairs $(s,a,s^{\prime})$ induced by policy $\pi$ under the dynamics $T$, defined by:
-$\rho_{T}^{\pi}(s,a,s^{\prime})=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}\mathbb{P}[s_{t}=s,a_{t}=a,\\ s_{t+1}=s^{\prime}|s_{0}\sim\mu_{0},a_{t}\sim\pi(\cdot|s_{t}),s_{t+1}\sim T(\cdot|s_{t},a_{t})].$
-Here, $\mu_{0}$ denotes the initial state distribution. For simplicity, we use shorthand notation $\rho^{\pi}_{T}$ in the remainder of the article.
-Offline Model-based RL.
-Offline model-based RL aims to find the optimal policy by leveraging a learned dynamics model $M$.
-Given the offline dataset $\mathcal{D}_{R}$, $M$ is typically trained using maximum likelihood estimation, formulated as $\min_{M}\mathbb{E}_{s,a,s^{\prime}\sim\mathcal{D}_{R}}[-\log M(s^{\prime}|s,a)]$.
-The training policy $\pi$ will undergo $k$-step rollout in learned dynamics model $M$ and the generated data will be added to a replay buffer $\mathcal{D}_{M}$.
-Then $\pi$ can be optimized using mini-batches of data sampled from $\mathcal{D}_{R}\cup\mathcal{D}_{M}$, where each datapoint is sampled from the offline dataset $\mathcal{D}_{R}$ with probability $f$ and the model dataset $\mathcal{D}_{M}$ with probability $1-f$.
-Throughout this work, we assume that the reward function $r(s,a,s^{\prime})$ is known, though it can also be treated as part of the model if unknown.
-4 Demystify OOD in Offline Model-based RL
 In this section, as illustrated in Fig.
-2, we explore the out-of-distribution (OOD) issues in offline model-based RL.
-We focus on the importance of aligning both the model and offline data, as well as the model and environment policy behavior, which motivates the direction of this research.
-4.1 OOD Issues in Offline RL
-In offline RL, OOD problems (Mao et al., 2024b) arise when agents encounter states or select actions that fall outside the distribution of the offline data during training or testing.
-OOD actions refer to actions that the behavior policy does not choose in specific states, rather than actions that are simply absent from the offline dataset.
-This implies that OOD actions are inherently state-dependent.
-Thus some studies use the term OOD state-action pairs to describe the state-action pair $(s,a)$ that does not appear in the offline dataset.
-Taking OOD actions during training can lead to severely inaccurate value estimation and significantly degraded policy performance in testing.
-Consequently, avoiding OOD actions has been a focus in previous offline RL research.
-OOD states primarily arise in three main scenarios: 1) The learned policy executes unreliable OOD actions, leading to transitions into OOD states.
-2) The initial state of the real environment lies outside the offline dataset.
-Additionally, stochastic dynamics could unexpectedly transition the agent into OOD states, even if it takes in-distribution (ID) actions in ID states.
-3) Nonstationary dynamics, such as disturbances in real-world environments, can introduce unexpected changes in state distributions.
-While this factor is outside the scope of our offline model-based setting, it is an important consideration in real-world applications.
-4.2 OOD Issues under Dynamics Model Integration
-OOD issues in offline model-based RL are intrinsically linked to the mismatch between synthetic and real data distributions, compounded in model-based settings, where model inaccuracies introduce additional sources of error.
-More extensive discussions are provided in Appendix B.
+2, we explore the OOD issues in offline MBRL.
+We focus on the essence of aligning both the synthetic and offline data, as well as the dynamics model and real environment policy behavior, which motivates the direction of this research.
+4.1. OOD Issues in Offline RL
+In offline RL, OOD problems Mao et al.
+(2024a) arise when agents encounter states or select actions that fall outside the offline data.
+Avoiding the OOD issue is a prerequisite for effective policy training.
+OOD actions refer to actions that the behavior policy does not choose in specific states, rather than actions that are absent from the offline dataset.
+The value of OOD actions is prone to overestimation due to limited sample coverage of the offline dataset and the requirement of maximizing the action values, leading to the incorrect or suboptimal direction of policy improvement.
+OOD states primarily arise in three main scenarios: 1) The policy executes unreliable OOD actions, leading to transitions into OOD states.
+2) Stochastic dynamics could transition the agent into OOD states, even if it takes in-distribution (ID) actions in ID states.
+Additionally, the initial state of the real environment may lie outside the dataset.
+3) Nonstationary dynamics, such as disturbances from real world Luo et al.
+(2024a) and cross-domain settings Lyu et al.
+(2025, 2024) can introduce fluctation in state distributions.
+In this paper, we consider the first two factors.
+While the third factor is outside the scope of the offline RL setting, it is a critical consideration in some real-world applications.
+4.2. OOD Issues under Biased Dynamics Model
+The prediction bias and stochasticity of the dynamics model compound the OOD issue in the offline setting.
+The prediction bias causes a misalignment between synthetic and offline data, and this bias is amplified with iterative use of the model, leading to the undesirable exploitation of OOD state-action pairs Levine et al.
+(2020).
+Meanwhile, the stochasticity elevates the likelihood of encountering OOD states, undermining the policy performance across both learned models and real environments.
+More discussions are given in Appendix B.
 OOD actions.
-The model, $M$, is used to generate synthetic data for training, but it is essential to avoid taking OOD actions when rolling out policy in the model.
-This can trigger negative model exploitation (Levine et al., 2020), resulting in poor policy performance.
-Aligning the distributions of synthetic and offline data can effectively mitigate these issues as it ensures accurate value estimation.
+The integration of the dynamics model allows the agent to query the possible outcomes of selecting arbitrary actions in any given state.
+However, the dynamics model cannot provide accurate predictions for OOD actions, and the prediction bias will mislead the estimation of these OOD action values.
+To address the problem, it is essential to constrain the scope of counterfactual queries within the support of the offline dataset, which amounts to aligning the distribution of synthetic data with the offline data.
 OOD states.
-The performance of the learning policy during deployment is critically dependent on the alignment of transition pairs between the learned model and the real environment.
-Significant discrepancies between these transitions can cause performance degradation and lead to OOD states, highlighting the necessity of correcting transition biases during training.
-Even if the learned policy performs well in the model, it is important, during real-environment deployment, to avoid OOD states, for which we cannot accurately estimate the value function in the offline setting.
-In summary, while optimizing the policy $\pi$ to maximize returns, we should constrain it in two directions: 1) aligning the synthetic data $(s,a,s^{\prime})_{M}^{\pi}$ with the offline data $(s,a,s^{\prime})_{T}^{\pi_{\beta}}$, and 2) ensuring that $\pi$ exhibits consistent behavior in both dynamics models and real environments.
-5 Dual Alignment Maximin Optimization
-Inspired by the analysis above, we propose Dual Alignment Maximin Optimization (DAMO), a unified approach to mitigate both OOD actions and states.
-We further explore the distinct roles of the inner minimization and the outer maximization in addressing the distribution shift challenge.
-5.1 Maximin Objective for Unified Shift Mitigation
-We start from the root cause of distribution shifts, i.e., the discrepancy between learned and behavior policy behaviors in the underlying environment.
-To tackle this, we introduce a regularized objective that constrains the difference between these behaviors while optimizing for expected rewards:
+The agent can mitigate the OOD states stemming from the first scenario by constraining the training policy to avoid OOD actions during training.
+However, the inherent prediction bias and intrinsic stochasticity of the biased dynamics model will increase the potential of encountering OOD states stemming from the second scenario during deployment.
+To address this problem, it is essential to ensure that the policy behavior remains consistent across both the dynamics model and the real environment.
+In summary, while optimizing the training policy $\pi$ to maximize the expected returns, we should constrain the policy in two directions: 1) aligning the synthetic data $(s,a,s^{\prime})_{M}^{\pi}$ with the offline data $(s,a,s^{\prime})_{T}^{\pi_{\beta}}$, and 2) ensuring that training policy $\pi$ exhibits consistent behaviors in both dynamics models $M$ and real environments $T$.
+5. Dual Alignment Maximin Optimization
+Inspired by the analysis above, we propose Dual Alignment Maximin Optimization (DAMO), a unified maximin approach to mitigate both OOD actions and states through dual alignment.
+We further explore the distinct roles of the inner minimization and the outer maximization of DAMO in addressing the distribution shift.
+5.1. Maximin Objective for Shift Mitigation
+We start from the root cause of distribution shifts, i.e., the discrepancy between learned and behavior policy behaviors in the underlying environment $T$.
+To tackle this, we imitate the behavior regularization methods and introduce a regularized objective that constrains the difference between distributions of learned and behavior policy while optimizing for expected rewards:
 $\max\limits_{\pi}\mathbb{E}_{\rho_{T}^{\pi}}[r(s,a,s^{\prime})]-\alpha D_{KL}(\rho_{T}^{\pi}\|\rho_{T}^{\pi_{\beta}}),$
-where $\alpha$ is a regularization hyperparameter. Next, we self-consistently incorporate synthetic data through a divergence upper bound, deriving the following surrogate objective:
+where $\alpha$ is a regularization hyperparameter.
+Next, we self-consistently incorporate synthetic data through establishing an upper bound of KL divergence, deriving the following surrogate objective:
 $\max\limits_{\pi}\mathbb{E}_{\rho_{T}^{\pi}}[r(s,a,s^{\prime})-\alpha\log\dfrac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}]-\alpha D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi}).$ (1)
-This surrogate objective integrates two essential elements: a data alignment regularizer $\log(\rho_{M}^{\pi}/\rho_{T}^{\pi_{\beta}})$ to ensure compatibility between synthetic and offline data, and a behavior alignment regularizer $D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})$ that maintains consistent policy performance across models and real environments.
-While this objective (1) theoretically addresses the synthetic-to-real distribution mismatch (as shown later in Sec.
-5.3), it is impractical to directly compute due to the inaccessibility of $\rho_{T}^{\pi}$.
-To make this objective tractable, we apply the convex conjugate theorem into the dual formulation of reinforcement learning, reformulating it as a maximin optimization problem.
-This reformulation forms the core of our proposed DAMO, a unified maximin framework for mitigating distribution shifts in offline model-based RL.
-Theorem 5.1.
+Where $f$ is a convex function that satisfies the inequality: $f(x)\geq x\log x,\forall x>0.$ This surrogate objective integrates two essential elements: a data alignment regularizer $\log(\rho_{M}^{\pi}/\rho_{T}^{\pi_{\beta}})$ to ensure compatibility between synthetic and offline data, and a behavior alignment regularizer $D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})$ that maintains consistent policy performance across dynamics models $M$ and real environments $T$.
+While objective (1) theoretically addresses the synthetic-to-real distribution mismatch (as shown later in Sec.
+5.3), it is impractical to compute due to the inaccessibility of distribution $\rho_{T}^{\pi}$.
+To address this problem, we apply the convex conjugate theorem to the dual formulation of the RL objective, reformulating it as a maximin optimization problem.
+This reformulation forms the core of DAMO, a unified maximin framework for mitigating the OOD issue.
+Theorem. 1.
 (1) is equivalent to the following problem:
-$\max\limits_{\pi}\min\limits_{\tilde{Q}(s,a)}(1-\gamma)\mathrm{E}_{s\sim\mu_{0},a\sim\pi}[\tilde{Q}(s,a)]\\ +\alpha\mathrm{E}_{\rho_{M}^{\pi}}[f_{\star}(\Phi(s,a,s^{\prime})/\alpha)],$ (2)
-where $f_{\star}$ is the conjugate function of the convex function $f$, and $\Phi(s,a,s^{\prime})=\log r(s,a,s^{\prime})-\alpha\log\frac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}+\tau^{\pi}Q(s,a)-Q(s,a)$, with $\tau^{\pi}Q(s,a)=\gamma\sum_{a^{\prime}}Q(s^{\prime},a^{\prime})\pi(a^{\prime}|s^{\prime})$.
-Here, $\mu_{0}$ denotes the distribution of the initial state, and $\alpha$ is a hyperparameter to control the degree of conservatism.
-5.2 Practical Implementation of DAMO
+$\max\limits_{\pi}\min\limits_{Q(s,a)}(1-\gamma)\mathbb{E}_{s\sim\mu_{0},a\sim\pi}[Q(s,a)]+\alpha\mathbb{E}_{\rho_{M}^{\pi}}[f_{\star}(\Phi(s,a,s^{\prime})/\alpha)],$ (2)
+where $f_{\star}(x)$ is the conjugate function of the convex function $f(x)$, and $\Phi(s,a,s^{\prime})=r-\alpha\log\frac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}+\gamma\sum\limits_{a^{\prime}}Q(s^{\prime},a^{\prime})\pi(a^{\prime}|s^{\prime})-Q(s,a)$
+The proof of Theorem 1 utilizes the DICE to correct the discrepancies between distributions and can be found in Appendix A.2.
+5.2. Practical Implementation of DAMO
 To implement the maximin optimization (2) practically, we employ a classifier to approximate the data alignment term $\log(\rho_{M}^{\pi}/\rho_{T}^{\pi_{\beta}})$.
-Specifically, we train a classifier $h(s,a,s^{\prime})$ using the following loss function to distinguish transitions sampled from offline data and those generated by the model:
+We train a classifier $h(s,a,s^{\prime})$ with the following loss function to distinguish transitions sampled from offline data and synthetic data:
 $\min_{h}\dfrac{1}{|\mathcal{D}_{R}|}\sum_{(s,a,s^{\prime})\in\mathcal{D}_{R}}\log h(s,a,s^{\prime})\\ +\dfrac{1}{|\mathcal{D}_{M}|}\sum_{(s,a,s^{\prime})\in\mathcal{D}_{M}}[\log(1-h(s,a,s^{\prime}))].$ (3)
-Here, $\mathcal{D}_{R}$ and $\mathcal{D}_{M}$ represent the offline and the model-generated data, respectively.
-Based on the learned classifier $h^{\star}(s,a,s^{\prime})$, the data alignment modification $\log(\rho_{M}^{\pi}/\rho_{T}^{\pi_{\beta}})$ can be computed using the following analytical expression:
+$\mathcal{D}_{R}$ and $\mathcal{D}_{M}$ represent the buffer storing offline and synthetic data, respectively.
+Based on the learned classifier $h^{\star}(s,a,s^{\prime})$, $\log(\rho_{M}^{\pi}/\rho_{T}^{\pi_{\beta}})$ can be computed using the following analytical expression:
 $\log\dfrac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}=\log\dfrac{h^{\star}(s,a,s^{\prime})}{1-h^{\star}(s,a,s^{\prime})}.$ (4)
-Note that (2) can be divided into two phases: an inner value estimation phase and an outer policy improvement phase, resembling the structure of the actor-critic framework.
-This structure is directly adapted to implement DAMO.
-The complete DAMO framework is outlined in Algorithm 1.
-Algorithm 1 DAMO
-0:  Offline dataset $\mathcal{D}_{R}$, dynamics model $M$, actor network $\pi_{\phi}$, critic network $Q_{\theta}$, classifier $h_{\psi}$.
-1:  Train the probabilistic dynamics model $M$ on $\mathcal{D}_{R}$
-2:  Initialize the model dataset buffer $\mathcal{D}_{M}\leftarrow\varnothing$
-3:  for each epoch do
-4:     Generate synthetic rollouts by model $M$
-5:     Add transition data in these rollouts to $\mathcal{D}_{M}$
-6:     Update classifier $h_{\psi}$ according to Eq. (3)
-7:     Calculate $\log(\rho_{M}^{\pi}/\rho_{T}^{\pi_{\beta}})$ according to Eq. (4)
-8:     Update critic $Q_{\theta}$ via minimizing Eq. (2)
-9:     Update actor $\pi_{\phi}$ via maximizing Eq. (2)
-10:  end for
-5.3 Theoretical Insights for DAMO
-The core structure of DAMO involves mapping the inner minimization in (2) to the estimation of a dual conservative value $\tilde{Q}(s,a)$, and the outer maximization to policy improvement based on this value.
-In this subsection, we present deeper theoretical insights into the workings of DAMO.
+Note that the maximin optimization problem (2) can be decomposed into two phases: the inner value estimation and the outer policy improvement, resembling the actor-critic framework.
+Thus, we adopt the actor-critic structure to facilitate the implementation of DAMO.
+The complete framework is listed in Appendix C.1.
+5.3. Theoretical Insights for DAMO
+The core structure of DAMO involves mapping the inner minimization in (2) to the estimation of a dual conservative value $Q(s,a)$, and the outer maximization to policy improvement based on the estimated conservative value.
+In this subsection, we present deeper theoretical insights into the working mechanism of DAMO.
 Dual Conservative Value Estimation.
-We first demonstrate the existence of both explicit and implicit reward penalties in the optimization objective (2), which contribute to the construction of the dual conservative value estimation.
-Theorem 5.2.
-The optimal solution $Q_{\pi}^{\star}$ for the inner minimization optimization in (2) satisfies the following equation:
-$Q_{\pi}^{\star}=r-\alpha\log\dfrac{\rho^{\pi}_{M}}{\rho^{\pi_{\beta}}_{T}}-\alpha f^{\prime}(\dfrac{\rho^{\pi}_{T}}{\rho^{\pi}_{M}})+\tau^{\pi}Q_{\pi}^{\star}.$
-This indicates that DAMO inherently introduces an explicit data alignment modification, $\log(\rho^{\pi}_{M}/\rho^{\pi{\beta}}_{T})$, and an implicit behavior alignment adjustment, $f^{\prime}(\rho^{\pi}_{T}/\rho^{\pi}_{M})$.
-Minimizing the inner objective effectively reshapes the reward function, $r(s,a,s^{\prime})$, into a refined reward, $\tilde{r}(s,a,s^{\prime})$, which represents the dual conservative value estimation:
+We first show that the inner minimization within (2) includes both: (1) an explicit penalty that directly aligns the synthetic data with offline data, and (2) an implicit penalty that emerges during minimization, ensuring policy behavior consistency as the solution converges to the optimum.
+Theorem. 2.
+The optimal solution $Q_{\pi}^{\star}$ for the inner minimization optimization in (2) satisfies:
+$Q_{\pi}^{\star}=r-\alpha\log\dfrac{\rho^{\pi}_{M}}{\rho^{\pi_{\beta}}_{T}}-\alpha f^{\prime}(\dfrac{\rho^{\pi}_{T}}{\rho^{\pi}_{M}})+\mathcal{T}^{\pi}Q_{\pi}^{\star}.$
+The proof of Theorem 2 is provided in Appendix A.3.
+This theorem indicates that DAMO inherently introduces an explicit data alignment regularizer, $\log(\rho^{\pi}_{M}/\rho^{\pi{\beta}}_{T})$, and an implicit behavior alignment adjustment, $f^{\prime}(\rho^{\pi}_{T}/\rho^{\pi}_{M})$.
+The inner minimization thus reshapes the vanilla reward $r(s,a,s^{\prime})$ into a refined reward $\tilde{r}(s,a,s^{\prime})$:
 $\tilde{r}=r-\alpha\log\frac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}-\alpha f^{\prime}(\dfrac{\rho^{\pi}_{T}}{\rho^{\pi}_{M}}).$
-The data alignment modification $\log(\rho^{\pi}_{M}/\rho^{\pi{\beta}}_{T})$ penalizes transition $(s,a,s^{\prime})$ that exhibit discrepancies between synthetic and offline data, promoting alignment with in-distribution transitions.
-This helps ensure accurate value estimation for these transitions.
+The data alignment regularizer $\log(\rho^{\pi}_{M}/\rho^{\pi{\beta}}_{T})$ penalizes transition $(s,a,s^{\prime})$ that exhibit discrepancies between synthetic and offline data, promoting alignment with in-distribution transitions.
+This alignment helps ensure accurate value estimation for these transitions.
 Meanwhile, the behavior alignment adjustment $f^{\prime}(\rho^{\pi}_{T}/\rho^{\pi}_{M})$ focuses on that different dynamics impact the distribution of $(s,a,s^{\prime})$.
-By penalizing these discrepancies, we encourage the agent to select policies that exhibit similar performance in both the underlying environment and the dynamics model, thus mitigating the impact of OOD states and ensuring consistent policy behavior between dynamics models and real environments.
+By penalizing these discrepancies, we encourage the agent to select policies that exhibit similar performance in both the dynamics model and the underlying environment, thus mitigating the impact of OOD states during deployment and ensuring consistent policy behavior between dynamics models and real environments.
 Consistent Policy Improvement.
-Unlike conventional offline model-based approaches that optimize policies purely by maximizing estimated values, DAMO ensures the consistency between value estimation and policy improvement.
-Specially, DAMO employs $\tilde{Q}(s,a)$ to minimize the objective (2) and utilizes $\pi$ to maximize it.
-Conventional approaches typically follow a bilevel optimization framework, while DAMO employs a maximin paradigm, ensuring consistency and preserving dual conservatism.
+Unlike conventional offline actor-critic frameworks that optimize policies solely by maximizing estimated values, DAMO ensures the objective consistency between value estimation and policy improvement.
+Specifically, DAMO employs the conservative Q-value to minimize the objective (2) and utilizes $\pi$ to maximize it.
+Prior actor-critic approaches typically follow a bilevel optimization framework, while DAMO employs a maximin paradigm, ensuring objective consistency.
 As demonstrated in Sec.
-6.1, this consistent structure maintains the conservatism of value estimation during policy improvement.
-Once the inner minimization in (2) is solved, the outer maximization problem is equivalent to solving the surrogate objective (1), which provides a lower bound for the standard RL objective $\mathcal{J}(\pi)=\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t},s_{t+1})]$:
-Theorem 5.3.
-(1) is a lower bound of $\mathcal{J}(\pi)$ for all $\pi$:
+6.1, this objective-consistency maintains the conservatism of value estimation during policy improvement.
+Once the inner minimization is solved, the outer maximization reduces to the surrogate objective (1), providing a lower bound for the standard RL objective $\mathcal{J}(\pi)=\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t},s_{t+1})]$:
+Theorem. 3.
+The surrogate objective (1) serves as a lower bound of $\mathcal{J}(\pi)$ for all $\pi$:
 $\mathcal{J}(\pi)\geq\mathbb{E}_{\rho_{T}^{\pi}}[r(s,a,s^{\prime})-\alpha\log\dfrac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}]-\alpha D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi}).$
-This demonstrates that the equivalence between objective (2) and the standard reinforcement learning objective ensures that the learned policy performs well in the environment.
-6 Experiments
-In this section, we focus on the following key aspects: 1) The capability of DAMO to handle synthetic-to-real distribution mismatch, as well as the contributions of its components.
-2) The comparative performance of DAMO against existing methods on standard offline RL benchmarks.
-3) The significant impact of hyperparameter settings and implementation approaches on the overall performance of DAMO.
-We delve into these aspects using the D4RL benchmark (Fu et al., 2020) on the MuJoCo simulator (Todorov et al., 2012).
-Our implementation is based on the OfflineRL-Kit library [^1] https://github.com/yihaosun1124/OfflineRL-Kit, a comprehensive and high-performance library for implementing offline reinforcement learning algorithms.
-The basic parameters of DAMO are consistent with the settings of this library.
-The details of each experiment and the specific hyperparameters configuration can be found in Appendix D.
-6.1 Effectiveness of DAMO Paradigm
-In this section, we present empirical evidence demonstrating the dual capabilities of DAMO: 1) effectively aligning model-generated data with offline data distributions, and 2) consistently maintaining policy stability between dynamics models and real environments.
+The detailed proof of Theorem 3 can be found in Appendix A.1.
+This theorem demonstrates the equivalence between maximin objective (2) and the standard RL objective, thereby offering a theoretical guarantee that the policy learned through objective (2) can achieve consistent performance in the real environment.
+6. Experiments
+In this section, we focus on the following aspects: 1) The capability of DAMO to handle distribution mismatch, and the contributions of each component of DAMO.
+2) The performance of DAMO against existing methods on standard benchmarks.
+3) The impact of hyperparameter settings and implementation approaches.
+4) The OOD generalization ability of DAMO in the task with relabeled data.
+We delve into these aspects using the D4RL Fu et al.
+(2020) and NeoRL Qin et al.
+(2022) benchmark on the MuJoCo simulator Todorov et al.
+(2012).
+Our implementation is based on the OfflineRL-Kit library [^1] https://github.com/yihaosun1124/OfflineRL-Kit.
+The basic parameters of DAMO are consistent with the settings of the library.
+The experiment details and the hyperparameter configuration is listed in Appendix E.
+(a) Normalized Score
+(b) Estimated Value
+Figure 3.
+Training process of DAMO, w/o ir, w/o er.
+(a) The data alignment term is the primary factor that influences policy improvement, while the absence of the behavior alignment term results in performance degradation in the final stages.
+(b) The lack of the data alignment term leads to the overestimation of the value.
+6.1. Effectiveness of DAMO Paradigm
+In this section, we present empirical evidence demonstrating the dual capabilities of DAMO: 1) effectively aligning synthetic data with offline data, and 2) consistently maintaining policy stability between the dynamics models and real environments.
 Our detailed ablation study thoroughly investigates two key components of DAMO: inner value estimation and outer policy improvement.
+6.1.1. Dual Conservative Value Estimation
+In this part, we present an ablation study examining the contributions of the data alignment regularizer and behavior alignment regularizer.
+We evaluate their impact by removing each term from DAMO and training policies on the hopper-medium-expert task.
+Specifically, we test DAMO under three distinct conditions: the configuration without the data alignment regularizer (w/o er), the configuration without the behavior alignment regularizer (w/o ir), and the complete DAMO framework.
+The empirical results are presented in Fig.
+3.
+To further provide intuitive insights, we present the visualization of distributions under each setting in Fig 4.
 (a) The distribution of w/o er
@@
 (c) The distribution of DAMO
-Figure 3: (a) (b) (c) illustrate the results of DAMO without the data alignment term (w/o er), DAMO without the behavior alignment term (w/o ir), DAMO.
-The left panels depict the distributions of state-action pair $(s,a)$ for offline and synthetic data.
-The right panels display the distributions of states for real and synthetic data.
-(a) Normalized Score
-(b) Estimated Value
-Figure 4: Training process of DAMO, w/o ir, w/o er.
-(a) The data alignment term is the primary factor that influences policy improvement, while the absence of the behavior alignment term results in performance degradation in the final stages.
-(b) The lack of the data alignment term leads to the overestimation of the value.
-6.1.1 Dual Conservative Value Estimation
-In this part, we present an ablation study examining the individual contributions of the data alignment and behavior alignment terms.
-We systematically evaluate their impact by removing each term from DAMO and training policies on the hopper-medium-expert-v2 benchmark.
-To provide intuitive insights, we first visualize the state-action distributions of the final trained policies across three distinct dataset categories.
-These categories are defined as follows:
-  • Offline Data: The transition pairs $(s,a,s^{\prime})$ are collected with behavior policy $\pi_{\beta}$ in dataset Hopper-medium-expert.
-  • Synthetic Data: The transition pairs are generated by conducting training policy $\pi$ within the dynamics model $M$.
-  • Real Data: The transition pairs are generated by conducting training policy $\pi$ within the real environment $T$.
-Effectiveness of Data Alignment.
-To evaluate synthetic and offline data compatibility during training, we analyze the state-action pair distributions for both synthetic data and offline data under three distinct conditions: the complete DAMO framework, the configuration without the behavior alignment term (w/o ir), and the configuration without the data alignment term (w/o er).
-The visualization results are presented in the left column of Fig.
-3.
-Building on these initial observations, we then examine the training dynamics of DAMO under three specific conditions to further investigate their significant impacts.
-The results are presented in Fig.
-4.
-A comparative analysis between Fig.
-3(a) and Fig.
-3(c) reveals that the w/o er configuration fails to establish proper alignment between synthetic data and offline data, thereby inadequately mitigating OOD state-action pairs.
-The results in Fig.
-3(b) further highlight that the data alignment term enables w/o ir configuration to align synthetic data with offline data.
-To clarify the detrimental effects of incompatibility between these two data sources, Fig.
-4(b) illustrates that w/o er overestimates the value of OOD actions, leading to suboptimal training performance.
-This is corroborated by the poor evaluation scores of w/o er depicted in Fig.
-4(a).
-Effectiveness of Behavior Alignment.
-To assess the model-environment consistency during the testing phase, we follow a similar evaluation procedure.
-We visualize the state distributions of real environment and model-generated data under the same conditions.
-The visualization results are presented in the right column of Fig.
-3.
-A comparative analysis between Fig.
-3(b) and Fig.
-3(c) reveals that while w/o ir successfully achieves synthetic-offline data compatibility and mitigates OOD actions, the absence of the behavior alignment term leads to incomplete mitigation of OOD states, manifesting policy inconsistency between dynamics model and real environment.
-Although the behavior alignment term imposes constraints, the results in Fig.
-3(a) demonstrate that w/o er fails to effectively mitigate OOD states.
-The observed discrepancies in state distributions between real environment data and model-generated data under the w/o er highlight the critical importance of establishing robust compatibility between synthetic and offline data.
-This compatibility is essential for ensuring consistent policy performance across dynamics models and real environments.
-The influence is reflected in Fig.4(a), which shows performance degradation of w/o ir in the final stage.
-6.1.2 Towards Consistent Policy Improvement
-Table 1: The average value of OOD state-action pairs, $\infty$ for severe overestimation. The smallest values are bolded, which indicates the best mitigation of the overestimation problem.
-halfcheetah hopper walker2d
-Consistent Version 95.2 213.0 231.7
-Inconsistent Version 101.4 219.4 $\infty$
-$Real\ value$ 819.6 269.6 299.7
-This section presents a statistical analysis of the value estimation performance of DAMO across multiple datasets, with comparisons against the true value function.
-Our results validate that the compatible policy improvement framework effectively maintains the conservatism of value estimation.
-To investigate the significance of objective alignment in DAMO, we conduct a controlled comparison experiment by directly integrating DAMO with standard off-policy mechanisms (Inconsistent Version) while evaluating the critic’s estimation of state-action pair values and comparing it with the original DAMO (Consistent Version).
-We perform this analysis on the medium-expert dataset across three distinct environments: HalfCheetah, Hopper, and Walker2d.
-The obtained quantitative results are summarized in Table 1.
-The experimental findings reveal that the Consistent Version achieves conservative value estimations across all environments.
-Although the Inconsistent version maintains conservatism in the HalfCheetah and Hopper environments, it exhibits significant overestimation in the Walker2d environment.
-These results collectively demonstrate the critical role of objective alignment in consistently ensuring conservative value estimation and preventing overestimation errors.
-Table 2: Results on the D4RL Gym benchmark.
+Figure 4.
+(a)-(c) illustrate the results of DAMO without the data alignment term (w/o er), DAMO without the behavior alignment term (w/o ir), and standard DAMO.
+Left panel: state-action (s, a) distributions (offline vs.
+synthetic).
+Right panel: state distributions $s$ (real vs.
+synthetic).
+Table 1.
+Results on the D4RL Gym benchmark.
 The numbers reported for DAMO are the normalized scores averaged over the final iteration of training across 4 seeds, with $\pm$ standard deviation.
-Since we haven’t obtained the results of O-DICE on random datasets, we didn’t include it in the table and used a dash (-) to indicate the absence of data.
+Since we haven’t obtained the results of O-DICE on random datasets, we didn’t include it in the table and used a dash (-) to indicate the absence of scores.
 The top three scores are highlighted in bold.
 We use an asterisk (*) for the best score, an underscore (_) for the second, and bold text for the third.
-Task Name Algae Opti O-DICE MOPO COMBO TT RAMBO MOBILE DAMO (Ours)
+Task Name Algae Opti O-DICE MOPO COMBO $T$ RAMBO MOBILE DAMO (Ours)
 halfcheetah-random -0.3 11.6 - 38.5 38.8 6.1 $\textbf{39.5}^{*}$ 38.3 $34.0\pm 3.0$
@@
 walker2d-medium-replay 0.6 21.6 83.6 85.9 56.0 82.6 89.2 80.5 89.4 $\pm$ $\textbf{3.0}^{*}$
-halfcheetah-medium-expert -0.8 91.1 93.2 83.6 90.0 95.0 95.4 101.4 103.1 $\pm$ 0.9
+halfcheetah-medium-expert -0.8 91.1 93.2 83.6 90.0 95.0 95.4 101.4 103.1 $\pm$ $\textbf{0.9}^{*}$
 hopper-medium-expert 1.1 111.5 110.8 74.6 111.0 110.0 88.2 107.7 112.3 $\pm$ $\textbf{3.6}^{*}$
@@
 average score 0.0 43.4 - 68.7 66.0 61.1 66.1 72.2 $\textbf{74.4}^{*}$
-6.2 Comparison Results
+Effectiveness of Data Alignment.
+A comparative analysis between Fig.
+4(a) and Fig.
+4(c) reveals that w/o er fails to establish alignment between synthetic and offline data, thereby inadequately mitigating OOD state-action pairs.
+The results in Fig.
+4(b) further highlight that the data alignment regularizer enables the w/o ir configuration to align synthetic data with offline data.
+To clarify the detrimental effects of incompatibility between these two data sources, the result in Fig.
+3(b) illustrates that the w/o er configuration overestimates the value of OOD actions, leading to suboptimal training performance.
+This is corroborated by the poor evaluation score of the w/o er configuration depicted in Fig.
+3(a).
+Effectiveness of Behavior Alignment.
+A comparison between Fig.
+4(b) and Fig.
+4(c) reveals that while w/o ir successfully achieves synthetic-offline data compatibility and mitigates OOD actions, the absence of the behavior alignment term leads to incomplete mitigation of OOD states, manifesting policy inconsistency between dynamics models and real environments.
+Although the behavior alignment term imposes constraints, the results in Fig.
+4(a) demonstrate that w/o er fails to effectively mitigate OOD states.
+The observed discrepancies in state distributions between real environment data and model-generated data under the w/o er highlight the critical importance of establishing robust compatibility between synthetic and offline data.
+This compatibility is essential for ensuring consistent policy performance across dynamics models and real environments.
+The influence is reflected in Fig.3(a), which shows performance degradation of w/o ir in the final stage of training.
+6.1.2. Towards Consistent Policy Improvement
+This section presents a statistical analysis of the value estimation performance of DAMO across multiple datasets, with comparisons against the true value function of the policy in the real environment.
+Our results validate that the compatible policy improvement framework effectively maintains the conservatism of value estimation.
+To investigate the significance of objective alignment in DAMO, we conduct a comparison experiment by directly integrating DAMO with standard off-policy mechanisms (Inconsistent Version) while evaluating the critic’s estimation of state-action pair values and comparing it with the original DAMO (Consistent Version).
+The detailed implementation of the Inconsistent Version can be found in Appendix C.3.
+We perform this analysis on the medium-expert dataset across three distinct environments: HalfCheetah, Hopper, and Walker2d.
+The obtained quantitative results are summarized in Table 2.
+The experimental findings reveal that the Consistent Version achieves conservative value estimations across all environments.
+Although the Inconsistent version maintains conservatism in the HalfCheetah and Hopper environments, it exhibits significant overestimation in the Walker2d environment.
+These results demonstrate the critical role of objective alignment in persisting conservative value estimation and preventing overestimation.
+Table 2. The average value of OOD state-action pairs, $\infty$ for severe overestimation. The smallest values are bolded, which indicates the best mitigation of the overestimation problem.
+halfcheetah hopper walker2d
+Consistent Version 95.2 213.0 231.7
+Inconsistent Version 101.4 219.4 $\infty$
+RealvalueReal\ value 819.6 269.6 299.7
+6.2. Comparison Results
+Table 3. Results on the NeoRL benchmark. We report the normalized scores averaged over the trained policy across 4 seeds. The results of MOBILE are obtained by executing the official implementation.
+Task Name BC CQL TD3+BC EDAC MOPO MOBILE DAMO (Ours)
+Halfcheetah-L 29.1 38.2 30.0 31.3 40.1 54.8 $47.55\pm 3.11$
+Hopper-L 15.1 16.0 15.8 18.3 6.2 18.2 24.05 $\pm$ 2.18
+Walker2d-L 28.5 44.7 43.0 40.2 11.6 23.7 $\textbf{46.59}\pm\textbf{9.18}$
+Halfcheetah-M 49.0 54.6 52.3 54.9 62.3 79.0 $69.04\pm 1.13$
+Hopper-M 51.3 64.5 70.3 44.9 1.0 43.4 $64.57\pm 9.17$
+Walker2d-M 48.7 57.3 58.6 57.6 39.9 60.1 $\textbf{66.49}\pm\textbf{2.28}$
+Halfcheetah-H 71.3 77.4 75.3 81.4 65.9 71.8 $67.23\pm 17.5$
+Hopper-H 43.1 76.6 75.3 52.5 11.5 42.3 $74.87\pm 6.39$
+Walker2d-H 72.6 75.3 69.6 75.5 18.0 71.9 $74.28\pm 2.03$
+6.2.1. D4RL
 Dataset.
-We evaluate DAMO’s performance using the D4RL benchmark (Fu et al., 2020) within the MuJoCo simulation environment.
-Our comprehensive assessment covers 12 distinct datasets across three environments (HalfCheetah, Hopper, and Walker2d) and four dataset categories: random, medium, medium-replay, and medium-expert.
-Following standard evaluation protocols, we exclusively utilize the ”v2” versions of all datasets for consistent comparison.
+We first evaluate the performance of DAMO using the D4RL Fu et al.
+(2020).
+Our assessments cover 12 distinct datasets across three control environments (HalfCheetah, Hopper, and Walker2d) and four dataset categories: random, medium, medium-replay, and medium-expert.
+Following standard evaluation protocols, we utilize the ”v2” versions of all datasets in D4RL for consistent comparison.
 Baseline.
-We compare DAMO with two types of offline RL baselines: DICE-based algorithms which follow a similar form of training objective with DAMO and Model-based algorithms which achieve SOTA performance.
-For DICE-based approaches, Algae-DICE (Nachum et al., 2019b) transforms the intractable state-action-level constraint into a unified objective for policy training.
-Opti-DICE (Lee et al., 2021) directly estimates the stationary distribution corrections of the optimal policy to attain a high-rewarding policy.
-O-DICE (Mao et al., 2024a) using the orthogonal-gradient update to diminish the confliction of different gradients during policy training.
-For model-based approaches, MOPO (Yu et al., 2020) penalizes rewards via the predicted variance of ensembled dynamics models.
-COMBO (Yu et al., 2021) implements CQL within a model-based framework.
-TT (Janner et al., 2019) uses a transformer to model offline trajectories and employs beam search for planning.
-RAMBO (Rigter et al., 2022) adversarially trains the policy and the dynamics model within a robust framework.
-MOBILE (Sun et al., 2023) penalizes the value learning of synthetic data based on the estimated uncertainty of the Bellman Q-function, derived from an ensemble of models.
+We compare DAMO with two types of offline RL baselines: DICE-based algorithms, which follow a similar form of training objective to DAMO, and Model-based algorithms, which achieve SOTA performance in D4RL benchmark.
+For DICE-based approaches, Algae-DICE Nachum et al.
+(2019b) transforms the intractable state-action-level constraint into a unified objective for policy training.
+Opti-DICE Lee et al.
+(2021) directly estimates the stationary distribution corrections of the optimal policy to attain a high-rewarding policy.
+O-DICE Mao et al.
+(2024b) uses the orthogonal gradient update to diminish the conflict of different gradients during policy training.
+For model-based approaches, MOPO Yu et al.
+(2020) penalizes rewards via the predicted variance of ensembled dynamics models.
+COMBO Yu et al.
+(2021) implements CQL within a model-based framework.
+TT Janner et al.
+(2019) uses a transformer to model offline trajectories and employs beam search for planning.
+RAMBO Rigter et al.
+(2022) adversarially trains the policy and the dynamics model within a robust framework.
+MOBILE Sun et al.
+(2023) penalizes the value learning of synthetic data based on the estimated uncertainty of the Bellman Q-function, which is derived from an ensemble of models.
 Comparison Results.
-Table 2 reports the scores in D4RL benchmark.
+Table 1 reports the scores in the D4RL benchmark.
 Overall, DAMO demonstrates superior performance, achieving the highest average score across all baseline methods.
-Notably, for high-quality datasets such as medium-expert and medium-replay, DAMO consistently attains state-of-the-art (SOTA) or near-SOTA performance.
-These results suggest that effective OOD constraints, combined with a high-quality behavior policy, enable the agent to learn high-reward policies within in-distribution (ID) regions.
+Notably, for high-quality datasets such as medium-expert and medium-replay, DAMO consistently attains SOTA or near-SOTA performance.
+These results suggest that effective dual alignment, combined with a high-quality behavior policy, enable the agent to learn high-reward policies within ID regions.
 Furthermore, DAMO maintains strong performance on medium-quality datasets.
 This demonstrates the robustness of DAMO in handling less-than-ideal data conditions and highlights the importance of ensuring both model-environment policy consistency and synthetic-offline data compatibility.
-However, the performance of DAMO on random datasets reveals limitations, indicating the need for more aggressive exploration strategies in low-quality data scenarios to identify regions beyond those covered by the offline dataset.
-This phenomenon is particularly pronounced in the Walker2d-random dataset, where DAMO struggles to learn an effective policy.
-One possible explanation is that the optimization landscape of the environment may contain numerous suboptimal solutions corresponding to saddle points in the maximin framework.
-If this is the case, such structural characteristics could potentially make DAMO more susceptible to becoming trapped in local optima, thereby limiting its ability to discover better solutions.
-However, further investigation is needed to confirm this hypothesis.
-6.3 Effects of Coefficient $\alpha$
+However, the performance of DAMO on random datasets reveals limitations to some extent, indicating the need for more aggressive exploration strategies in low-quality data scenarios to identify uncovered regions.
+This phenomenon is pronounced in the Walker2d-random dataset, where DAMO struggles to learn an effective policy.
+One possible explanation is that the optimization landscape of the Walker2d contains numerous suboptimal solutions corresponding to saddle points in the maximin problem.
+If this is the case, such structural characteristics could potentially trap DAMO in local optima, limiting its ability to discover better solutions.
+Further investigation is needed to confirm the above hypothesis.
+6.2.2. NeoRL
+Table 4. Average returns on halfcheetah-jump. Batch Mean and Batch Max denote the mean and maximum average return across all trajectories in the offline dataset, respectively.
+Environment Batch Mean Batch Max DAMO MOPO COMBO BEAR BRAC-v TD3+BC
+halfcheetah-jump -1184 2887 $4871\pm 416$ $5030\pm 435$ $4868\pm 175$ $186\pm 231$ $1038\pm 179$ $1340\pm 303$
+Dataset.
+To further assess the performance of DAMO in real-world scenarios, we test it in the NeoRL Qin et al.
+(2022).
+By emulating real-world data-collection scenarios, the datasets in NeoRL characterize a narrow scope and limited coverage, which impose significant challenges for offline RL.
+Our assessments include nine datasets across three environments (HalfCheetah, Hopper, Walker2d) and three dataset quality levels (L, M, H), corresponding to low, medium, and high.
+Baseline.
+We have not chosen the DICE-based method for comparison, since the results of these methods can not be found in the original paper or the NeoRL paper, and searching for optimal hyperparameters for each method would take excessive time.
+Instead, we select SOTA offline model-free methods to ensure the diversity of the baseline.
+For model-based approaches, we keep two representative methods, MOPO Yu et al.
+(2020) and MOBILE Sun et al.
+(2023) for comparison.
+For model-free approaches, BC Bain and Sammut (1995) imitates the behavior policy from the offline dataset.
+CQL Kumar et al.
+(2020) penalizes the Q-values of OOD samples by incorporating regularization into the value loss function.
+TD3+BC Fujimoto and Gu (2021) introduces a behavior cloning regularization into the TD3 objective.
+EDAC An et al.
+(2021) penalizes the Q-function based on uncertainty derived from an ensemble of Q-networks.
+Comparison Results.
+Table 3 reports the scores in the NeoRL benchmark.
+DAMO maintains competitive performance under more realistic and difficult scenarios, indicating strong potential for real-world applicability.
+Specifically, the strong generalization ability in narrow data coverage induced by suboptimal policy facilitates DAMO in achieving SOTA or near-SOTA performance on datasets of low-quality and medium-quality.
+Among high-quality datasets, DAMO secures a near-SOTA score in the Walker2d-H dataset and a high ranking score in the Hopper-H dataset.
+However, while training with the Halfcheetah-H dataset, DAMO fails to learn the optimal in-distribution policy, which contradicts our analysis in D4RL results.
+Through inspecting the training process, we observe that although DAMO achieves rapid policy improvement during the early phase, the subsequent training process exhibits significant fluctuations, with occasional steep drops in performance.
+We attribute this instability to limited data coverage caused by the near-deterministic nature of the policy and the noise in the environment, which together present a substantial challenge for robust learning.
+6.3. OOD generalization of DAMO
+Model-based methods are generally more capable of performing OOD generalization than model-free methods.
+As DAMO follows the idea of behavior regularization, which potentially hinders the agent from exploration, it remains unclear whether DAMO preserves the superior generalization ability Lyu et al.
+(2022); Liu et al.
+(2024).
+To clarify this problem, we evaluate DAMO in the halfcheetah-jump task Yu et al.
+(2020) where the agent solves a task that entails the effective generalization outside the scope of the offline data.
+Specifically, a behavior policy is trained online to control the cheetah to move forward as fast as possible.
+The rewards of the collected trajectories are then reshaped to encourage the cheetah jumping during offline training.
+Since the behavior policy is trained for a different purpose, the dataset only contains the suboptimal policy for jumping.
+Therefore, OOD generalization is crucial for the agent to acquire a high-return policy.
+We compare DAMO with other model-based methods and behavior-regularization methods.
+The implementation of the experiment and introduction of baselines are presented in Appendix C.5.
+As shown in Table 4, DAMO achieves comparable performance with model-based methods and outperforms all behavior-regularization methods.
+This suggests that combining MBRL with behavior regularization enables the agent to achieve both offline conservatism and OOD generalization when handling OOD issues in offline RL.
 (a) Tuning of $\alpha$
 (b) DAMO without fixing $\alpha$
-Figure 5: Two sets of experiments on hyperparameter $\alpha$.
-(a) The tuning experiment of $\alpha$ was performed on the hopper-medium-expert dataset.
+Figure 5.
+(a) The tuning experiment of $\alpha$ was performed on the hopper-medium-expert dataset for validation.
 (b) The fixing experiment of $\alpha$ was performed on the halfcheetah-medium-expert dataset for validation.
-DAMO is regulated by a key hyperparameter: the hyperparameter $\alpha$ which controls the extent of policy discrepancies.
-In practical implementation, we fix $\alpha$ in the actor training.
+6.4. Effects of Coefficient $\alpha$
+DAMO is regulated by a key hyperparameter: the hyperparameter $\alpha$, which controls the extent of policy discrepancies.
 While we consider $\alpha$ as a hyperparameter controlling the degree of conservatism, the empirical results in Fig.
 5(a) demonstrate that DAMO maintains robust performance across a wide range of $\alpha$ values, with its primary impact observed in the final stage performance.
-Notably, DAMO achieves superior performance with larger $\alpha$ values (e.g., $\alpha=3.0,5.0$), whereas insufficiently small values (e.g., $\alpha=1.0$) tend to promote excessive risk-taking during exploration, ultimately failing to converge to an optimal policy.
-To enhance training stability, we initially fix $\alpha$ to 1.0 in the actor training.
+DAMO achieves superior performance with larger $\alpha$ values (e.g., $\alpha=3.0,5.0$), whereas insufficiently small values (e.g., $\alpha=1.0$) tend to promote excessive risk-taking during exploration, failing to converge to an optimal policy.
+To enhance training stability, we initially follow the prior work Luo et al.
+(2024a) and fix $\alpha$ to 1.0 in the actor training.
+We present the concrete implementation in Appendix C.4.
 However, experimental results demonstrate that DAMO achieves competitive performance even without this specific implementation.
-To further validate this observation, we conduct comparative experiments on the HalfCheetah environment, evaluating both the standard DAMO implementation and its variant without fixed $\alpha$ under their respective optimal hyperparameter configurations.
+To further validate this observation, a comparative experiment is conducted, evaluating both the standard DAMO implementation and its variant without fixed $\alpha$ under their optimal hyperparameter configurations.
 As shown in Fig.
-5(b), both settings achieve a competitive performance, confirming that DAMO maintains robust performance regardless of the implementation of fixing $\alpha$.
-7 Conclusion and Discussion
-This paper critically examines limitations in existing offline model-based RL approaches, highlighting the crucial need to address behavior inconsistency caused by policy discrepancies between biased models and real environments.
-Building upon our analysis, we introduce a unified framework that concurrently ensures model-environment policy consistency and synthetic-offline data compatibility, with comprehensive experimental validation demonstrating its effectiveness.
+5(b), both settings achieve competitive performances, confirming that DAMO maintains robust performance regardless of the specific implementation.
+7. Conclusion and Discussion
+This paper examines limitations in existing offline MBRL approaches, highlighting the crucial need to address behavior inconsistency caused by policy discrepancies between biased models and real environments.
+Building upon our analysis, we introduce a unified framework that concurrently ensures both synthetic-offline data compatibility and model-environment policy consistency, with experimental validation demonstrating its effectiveness.
 Significantly, the maximin optimization framework of DAMO demonstrates the benefits of aligning the actor and critic training objectives within the actor-critic architecture, ensuring essential conservatism during policy updates.
-The limitations of our work are focused on two-fold: the sampling of the initial state and the computing of the data alignment term.
-Although we assume the distribution of the initial state is known, the offline dataset does not explicitly label which states are initial states.
-Therefore, we randomly sample a state from the offline dataset to serve as the initial state, which creates a gap between theory and practice.
-When calculating the data alignment term, we trained a classifier to assist in the process.
-However, the classifier’s accuracy is critical: imprecise classifiers fail to estimate the modified reward accurately, leading to unstable training and potentially invalidating the practical objective as a lower bound.
-Future work could extend DAMO to areas such as offline-to-online RL (Li et al., 2023), where methods with strong adaptability and the ability to bridge the synthetic-to-real distribution gap are in high demand.
-Impact Statement
-This work presents an effective solution for addressing the distribution shift in offline model-based RL through dual policy-environment alignment, enabling robust simulation-to-reality transfer in safety-critical domains such as medical robotics.
-While advancing certifiable policy deployment via implicit behavior constraints and adaptive value estimation, challenges remain in reward specification sensitivity and residual model bias mitigation.
-The framework highlights the critical need for continuous dynamic alignment calibration to ensure reliable real-world RL system performance.
+Our future work could extend DAMO to areas such as offline-to-online RL Li et al.
+(2023), where the method with adaptability and the capability to bridge the synthetic-to-real distribution gap is in great demand.
+The limitations of DAMO are focused on two-fold: the sampling of the initial state and the computation of the data alignment term.
+Although we assume the distribution of the initial state is accessible, some offline datasets do not explicitly label the initial states.
+Therefore, we randomly sample a state from these offline datasets to serve as the initial state, which creates a deployment gap.
+A classifier is employed to assist computation.
+However, the accuracy of the classifier is critical: imprecise classifiers fail to estimate the modified reward, leading to unstable policy training and potentially invalidating the objective as a lower bound.
 Appendix A Proofs in the Main Text
-We present a detailed proof process for each theorem discussed in the paper.
-The sequence of proofs follows the logic of derivation rather than strictly adhering to their order of appearance in the main text.
-Theorem A.1.
-The KL-divergence has the following upper bound:
-$D_{KL}(\rho_{T}^{\pi}\|\rho_{T}^{\pi_{\beta}})\leq\mathrm{E}_{\rho_{T}^{\pi}}[\log\dfrac{\rho_{T}^{\pi}}{\rho_{T}^{\pi_{\beta}}}]+D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi}).$
+We present a detailed proof for each theorem discussed above. The sequence of proofs follows the logic of derivation rather than strictly adhering to their order of appearance in the main text.
+A.1. Proof of Theorem 3
+To prove Theorem 3, we first introduce an upper bound of the KL-divergence.
+Lemma. 0.
+For any convex function $f(x)$ satisfying the inequality: $f(x)\geq x\log x,\forall x>0$, the KL-divergence admits the following upper bound:
+$D_{KL}(\rho_{T}^{\pi}\|\rho_{T}^{\pi_{\beta}})\leq\mathbb{E}_{\rho_{T}^{\pi}}[\log\dfrac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}]+D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi}).$
 Proof.
-$D_{KL}(\rho_{T}^{\pi}\|\rho_{T}^{\pi_{\beta}})$ $=\mathrm{E}_{\rho_{T}^{\pi}}[\log\dfrac{\rho_{T}^{\pi}}{\rho_{T}^{\pi_{\beta}}}]$
-$=\mathrm{E}_{\rho_{T}^{\pi}}[\log\dfrac{\rho_{T}^{\pi}}{\rho_{T}^{\pi_{\beta}}}\cdot\dfrac{\rho_{M}^{\pi}}{\rho_{M}^{\pi}}]$
-$=\mathrm{E}_{\rho_{T}^{\pi}}[\log\dfrac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}]+D_{KL}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})$
-$\leq\mathrm{E}_{\rho_{T}^{\pi}}[\log\dfrac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}]+D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})$
-Corollary A.2.
-We have the following surrogate objective
-$\arg\max\limits_{\pi}\mathrm{E}_{\rho_{T}^{\pi}}[r(s,a,s^{\prime})-\log\dfrac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}]-D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})$
-A discounted factor $\alpha$ will be added for training stability.
-$\arg\max\limits_{\pi}\mathrm{E}_{\rho_{T}^{\pi}}[r(s,a,s^{\prime})-\alpha\log\dfrac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}]-\alpha D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})$ (5)
-Lemma A.3.
-We can represent f-divergence in a function form: $D_{f}(x\|p)=\mathbb{E}_{z\sim p}f(\frac{x}{p})$, then its Fenchel conjugate function is $\mathbb{E}_{z\sim p}f_{\star}(y(z))$
+$D_{KL}(\rho_{T}^{\pi}\|\rho_{T}^{\pi_{\beta}})$ $=\mathbb{E}_{\rho_{T}^{\pi}}[\log\dfrac{\rho_{T}^{\pi}}{\rho_{T}^{\pi_{\beta}}}]$
+$=\mathbb{E}_{\rho_{T}^{\pi}}[\log\dfrac{\rho_{T}^{\pi}}{\rho_{T}^{\pi_{\beta}}}\cdot\dfrac{\rho_{M}^{\pi}}{\rho_{M}^{\pi}}]$
+$=\mathbb{E}_{\rho_{T}^{\pi}}[\log\dfrac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}]+D_{KL}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})$
+$\leq\mathbb{E}_{\rho_{T}^{\pi}}[\log\dfrac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}]+D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})$
+To establish the final inequality, we invoke the property that the f-divergence dominates the KL divergence, i.e., $D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})\geq D_{KL}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})$, which stems from the inequality $f(x)\geq x\log x$ imposed on the convex function $f(x)$.
+∎
+Deploying the Lemma 4 into the objective $\max\limits_{\pi}\mathbb{E}_{\rho_{T}^{\pi}}[r(s,a,s^{\prime})]-\alpha D_{KL}(\rho_{T}^{\pi}\|\rho_{T}^{\pi_{\beta}})$, then we have a lower bound of vanilla RL objective $\mathcal{J}(\pi)=\mathbb{E}_{\rho_{T}^{\pi}}[r(s,a,s^{\prime})]$.
+Corollary. 0.
+We have the following surrogate objective that serves as a lower bound of $\mathcal{J}(\pi)$:
+$\mathcal{J}(\pi)\geq\mathbb{E}_{\rho_{T}^{\pi}}[r(s,a,s^{\prime})-\alpha\log\dfrac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}]-\alpha D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi}).\quad\forall\alpha\geq 0$
 Proof.
-We use $g_{\star}(x)$ to denote $D_{f}(x\|p)$
-$g_{\star}(x)$ $=\sup\limits_{y\in\mathcal{X}}\{\langle x,y\rangle-D_{f}(y\|p)\}$
-[…]
-[…]
-$=\int_{R}p(z)\sup\limits_{z}[x(z)\frac{y(z)}{p(z)}-f(\frac{y(z)}{p(z)})]dz$
-$=\int_{R}p(z)\sup\limits_{z}[x(z)y(z)-f(y(z))]dz$
-[…]
-$=\mathrm{E}_{z\sim p}f_{\star}(y(z))$
-Corollary A.4.
-For f-divergence $D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})$
-$D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})=\max\limits_{y(s,a,s^{\prime})}\mathrm{E}_{\rho_{T}^{\pi}}[y(s,a,s^{\prime})]-\mathrm{E}_{\rho_{M}^{\pi}}[f_{\star}(y(s,a,s^{\prime}))]$
+The following objective is a lower bound of $\mathcal{J}(\pi)$ due to the non-negativity of KL-divergence:
+$\mathbb{E}_{\rho_{T}^{\pi}}[r(s,a,s^{\prime})]-\alpha D_{KL}(\rho_{T}^{\pi}\|\rho_{T}^{\pi_{\beta}})$
+We substitute the $D_{KL}(\rho_{T}^{\pi}\|\rho_{T}^{\pi_{\beta}})$ with its upper bound in theorem 4, then for any policy $\pi$, we have:
+$\mathcal{J}(\pi)\geq\mathbb{E}_{\rho_{T}^{\pi}}[r(s,a,s^{\prime})-\alpha\log\dfrac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}]-\alpha D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi}).$
+We have now completed the proof of Theorem 3.
+A.2. Proof of Theorem 1
+Next, we present the proof of Theorem 1. We first introduce two lemmas.
+Firstly, we define the conjugate functional and provide a corresponding proposition of biconjugate.
+Lemma. 0.
+Given a functional $f:\mathbb{E}\rightarrow\mathbb{R}$, $\mathbb{E}$ is a Hilbert space. Its Fenchel conjugate functional $f_{\star}:\mathbb{E}\rightarrow\mathbb{R}$ is defined as:
+$f_{\star}(y)=\sup\limits_{x\in\mathbb{E}}\{\langle y,x\rangle-f(x)\}$
+$\langle y,x\rangle$ is the inner dot of $y,x$.
+We call the functional $f_{\star\star}=(f_{\star})_{\star}$ as the biconjugate functional of $f$.
+If $f$ is a convex, closed functional, then its biconjugate functional is itself:
+$f_{\star\star}=f$
 Proof.
-Verifying $D_{f}$ is a convex, closed function. A convex and closed function’s double conjugate function is itself, so $D_{f}$ can be represented in a MAX form. ∎
-Theorem A.5.
-Surrogate objective (5) is equivalent to the following maximin problem
-$\max\limits_{\pi}\min\limits_{y(s,a,s^{\prime})}\mathrm{E}_{\rho_{T}^{\pi}}[r(s,a,s^{\prime})-\alpha\log\dfrac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}-\alpha y(s,a,s^{\prime})]+\alpha\mathrm{E}_{\rho_{M}^{\pi}}[f_{\star}(y(s,a,s^{\prime}))]$
+The proof of the proposition about the biconjugate functional can be found in Fenchel (2014)
+Then, in the second lemma, we consider the convex functional.
+Lemma. 0.
+We can represent f-divergence in a functional form with $x(z)$: $D_{f}(x\|p)=\mathbb{E}_{z\sim p}f(\frac{x(z)}{p(z)})\triangleq h_{p}(x)$, that means we fix the distributional function $p$ and treat $D_{f}(x\|p)$ as a functional $h_{p}(x)$.
+Then its Fenchel conjugate function is $\mathbb{E}_{z\sim p}f_{\star}(y(z))$
 Proof.
-According to lemma A.4, the objective can be presented in a Min form. ∎
-Theorem A.6.
-The following objective is equivalent to the surrogate objective
-$\max\limits_{\pi}\min\limits_{Q(s,a)}J(\pi,Q(s,a))=\max\limits_{\pi}\min\limits_{Q(s,a)}(1-\gamma)\mathrm{E}_{s\sim\mu_{0},a\sim\pi}[Q(s,a)]+\mathrm{E}_{\rho_{M}^{\pi}}f_{\star}(\Phi(s,a,s^{\prime})/\alpha)$ (6)
+We use $g_{\star}(y)$ to denote the conjugate functional of $D_{f}(x\|p)$, $\mathcal{X}$ to denote the $L^{2}space$. Then we have:
+$g_{\star}(y)$ $=\sup\limits_{x\in\mathcal{X}}\{\langle y,x\rangle-h_{p}(x)\}$
+$=\sup\limits_{x\in\mathcal{X}}\{\int_{R}y(z)x(z)dz-\int_{R}p(z)\cdot f(\frac{x(z)}{p(z)})dz\}$
+$=\sup\limits_{x\in\mathcal{X}}\{\int_{R}p(z)[y(z)\frac{x(z)}{p(z)}-f(\frac{x(z)}{p(z)})]dz\}$
+$=\int_{R}p(z)\sup\limits_{x(z)}[y(z)\frac{x(z)}{p(z)}-f(\frac{x(z)}{p(z)})]dz$
+$=\int_{R}p(z)\sup\limits_{x^{\prime}(z)}[y(z)x^{\prime}(z)-f(x^{\prime}(z))]dz$
+$=\int_{R}p(z)f_{\star}(y(z))dz$
+$=\mathbb{E}_{z\sim p}f_{\star}(y(z))$
+The first equality is the definition of the Fenchel conjugate function.
+In the second equality, we give the explicit expressions for the inner product and the KL divergence.
+In the fifth equality, we make the variable change and substitute $(x(z)/p(z))$ with $x^{\prime}(z)$.
+The sixth equality follows the definition of the Fenchel conjugate function.
+∎
+We can employ Lemma 6, 7 to represent the f-divergence in surrogate objective (1) in a MAX-form.
+Corollary. 0.
+The f-divergence $D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})$ satisfy
+$D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})=\max\limits_{y(s,a,s^{\prime})}\mathbb{E}_{\rho_{T}^{\pi}}[y(s,a,s^{\prime})]-\mathbb{E}_{\rho_{M}^{\pi}}[f_{\star}(y(s,a,s^{\prime}))].$
 Proof.
-Defining $y(s,a,s^{\prime})=\dfrac{r(s,a,s^{\prime})-\alpha\log\frac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}+\tau^{\pi}Q(s,a)-Q(s,a)}{\alpha}$.
+We can verify that $D_{f}(x\|p)$ is a convex, closed functional.
+Then the biconjugate functional of $D_{f}(x\|p)$ is itself according to Lemma 6.
+We keep the notation and use $g_{\star}(y)$ to denote the conjugate function of $D_{f}(x\|p)$, $y(s,a,s^{\prime})$ is any function defined in the transition space.
+Then $D_{f}(x\|p)$ is the conjugate of $g_{\star}(y)$.
+Therefore, $D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})$ can be represented by the conjugate functional of $g_{\star}(y)$:
+$D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})$ $=\max\limits_{y}\langle y,\rho^{\pi}_{T}\rangle-g_{\star}(y)$
+$=\max\limits_{y}\int\rho^{\pi}_{T}\cdot y\ \mathrm{d}s\ \mathrm{d}a\ \mathrm{d}s^{\prime}-g_{\star}(y)$
+$=\max\limits_{y}\mathbb{E}_{\rho_{T}^{\pi}}[y]-\mathbb{E}_{\rho_{M}^{\pi}}[f_{\star}(y)]$
+The first equality is the definition of the Fenchel conjugate function.
+In the second equality, we express the inner product in the form of an integral.
+In the third equality, we represent $g_{\star}(y)$ in the form of expectation according to Lemma 7.
+∎
+With the Corollary 8, we can attain a maximin form of surrogate objective (1).
+Theorem. 9.
+Optimizing surrogate objective (1) is equivalent to solving the following maximin problem
+$\max\limits_{\pi}\min\limits_{y}\mathbb{E}_{\rho_{T}^{\pi}}[r(s,a,s^{\prime})-\alpha\log\dfrac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}-\alpha y]+\alpha\mathbb{E}_{\rho_{M}^{\pi}}[f_{\star}(y)]$ (5)
+Proof.
+Substituting the $D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})$ in surrogate objective (1) with its MAX-form in Corollary 8, then we attain the equivalent maximin form of surrogate objective (1). ∎
+Finally, we should make the maximin objective feasible in the offline RL setting, which means that we should eliminate the expectation over $\rho^{\pi}_{T}$ in (5).
+Theorem. 10.
+The maximin objective (5) is equivalent to the following objective
+$\max\limits_{\pi}\min\limits_{Q(s,a)}(1-\gamma)\mathbb{E}_{s\sim\mu_{0},a\sim\pi}[Q(s,a)]+\alpha\mathbb{E}_{\rho_{M}^{\pi}}f_{\star}(\Phi(s,a,s^{\prime})/\alpha)$ (6)
+Proof.
+Defining $y(s,a,s^{\prime})=\frac{1}{\alpha}r(s,a,s^{\prime})-\alpha\log\frac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}+\gamma\mathbb{E}_{a^{\prime}\sim\pi}Q(s^{\prime},a^{\prime})-Q(s,a)$.
 For the convenience, we define $\beta_{t}(s,a,s^{\prime})$
 $\beta_{t}(s,a,s^{\prime})=Pr[s_{t}=s,a_{t}=a,s_{t+1}=s^{\prime}|s_{0}\sim\mu_{0},a_{t}\sim\pi,s_{t+1}\sim T]$
-Then, $\rho_{T}^{\pi}(s,a,s^{\prime})=(1-\gamma)\sum\limits_{t=0}^{\infty}\gamma^{t}\beta_{t}$. Similarly, we can define $\beta_{t}(s,a,s^{\prime},a^{\prime})$.
-Substituting $y(s,a,s^{\prime})$ with $[r(s,a,s^{\prime})-\alpha\log\frac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}+\tau^{\pi}Q(s,a)-Q(s,a)]/\alpha$ and we have:
-$\max\limits_{\pi}\min\limits_{Q(s,a)}$ $\mathrm{E}_{\rho_{T}^{\pi}}[Q(s,a)-\tau^{\pi}Q(s,a)]$
+According to definition, $\rho_{T}^{\pi}(s,a,s^{\prime})=(1-\gamma)\sum\limits_{t=0}^{\infty}\gamma^{t}\beta_{t}$. Similarly, we can define $\beta_{t}(s,a)$ and $\beta_{t}(s,a,s^{\prime},a^{\prime})$.
+$\beta_{t}(s,a)=Pr[s_{t}=s,a_{t}=a|s_{0}\sim\mu_{0},a_{t}\sim\pi,s_{t}\sim T]$
+$\beta_{t}(s,a,s^{\prime},a^{\prime})=Pr[s_{t}=s,a_{t}=a,s_{t+1}=s^{\prime},a_{t+1}=a^{\prime}|\mu_{0},\pi,T]$
+Substituting $y(s,a,s^{\prime})$ with $r(s,a,s^{\prime})-\alpha\log\frac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}+\gamma\mathbb{E}_{a^{\prime}\sim\pi}Q(s^{\prime},a^{\prime})-Q(s,a)]/\alpha$, then the first term in objective (5) can be represented as:
+$\mathbb{E}_{\rho_{T}^{\pi}}[r(s,a,s^{\prime})-\alpha\log\dfrac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}-\alpha y(s,a,s^{\prime})]$
+$=$ $\mathbb{E}_{\rho_{T}^{\pi}}[Q(s,a)-\gamma\mathbb{E}_{a^{\prime}\sim\pi}Q(s^{\prime},a^{\prime})]$
 $=$ $\sum\limits_{s,a,s^{\prime}}\rho_{T}^{\pi}(s,a,s^{\prime})[Q(s,a)-\gamma\sum\limits_{a^{\prime}}Q(s^{\prime},a^{\prime})\pi(a^{\prime}|s^{\prime})]$
@@
 $=$ $(1-\gamma)[\sum\limits_{s,a,s^{\prime}}\sum\limits_{t=0}^{\infty}\gamma^{t}\beta_{t}(s,a,s^{\prime})Q(s,a)-\sum\limits_{s,a,s^{\prime},a^{\prime}}\sum\limits_{t=0}^{\infty}\gamma^{t+1}\beta_{t}(s,a,s^{\prime})Q(s^{\prime},a^{\prime})\pi(a^{\prime}|s^{\prime})]$
-$=$ $(1-\gamma)[\sum\limits_{s,a,s^{\prime}}\sum\limits_{t=0}^{\infty}\gamma^{t}\beta_{t}(s,a,s^{\prime})Q(s,a)-\sum\limits_{s,a,s^{\prime},a^{\prime}}\sum\limits_{t=0}^{\infty}\gamma^{t+1}\beta_{t}(s,a,s^{\prime},a^{\prime})Q(s^{\prime},a^{\prime})]$
-$=$ $(1-\gamma)[\sum\limits_{s,a}\sum\limits_{t=0}^{\infty}\gamma^{t}\beta_{t}(s,a)Q(s,a)-\sum\limits_{s^{\prime},a^{\prime}}\sum\limits_{t=0}^{\infty}\gamma^{t+1}\beta_{t}(s^{\prime},a^{\prime})Q(s^{\prime},a^{\prime})]$
+$=$ $(1-\gamma)[\sum\limits_{s,a,s^{\prime}}\sum\limits_{t=0}^{\infty}\gamma^{t}\beta_{t}(s,a,s^{\prime})Q-\sum\limits_{s,a,s^{\prime},a^{\prime}}\sum\limits_{t=0}^{\infty}\gamma^{t+1}\beta_{t}(s,a,s^{\prime},a^{\prime})Q(s^{\prime},a^{\prime})]$
+$=$ $(1-\gamma)[\sum\limits_{s,a}\sum\limits_{t=0}^{\infty}\gamma^{t}\beta_{t}(s,a)Q-\sum\limits_{s^{\prime},a^{\prime}}\sum\limits_{t=0}^{\infty}\gamma^{t+1}\beta_{t}(s^{\prime},a^{\prime})Q(s^{\prime},a^{\prime})]$
 $=$ $(1-\gamma)\sum\limits_{s,a}\beta_{0}(s,a)Q(s,a)$
-$=$ $(1-\gamma)\mathrm{E}_{s\sim\mu_{0},a\sim\pi}[Q(s,a)]$
-Theorem A.7.
-The optimal solution of maximin question (6)’s inner question $Q^{\star}(s,a)$ satisfies
-$Q^{\star}(s,a)=r(s,a,s^{\prime})-\alpha\log\dfrac{\rho^{\pi}_{M}}{\rho^{\pi_{\beta}}_{T}}-\alpha f^{\prime}(\dfrac{\rho^{\pi}_{T}}{\rho^{\pi}_{M}})+\tau^{\pi}Q^{\star}(s,a)$ (7)
+$=$ $(1-\gamma)\mathbb{E}_{s\sim\mu_{0},a\sim\pi}[Q(s,a)]$
+In the first equality, we make a variable change of $y(s,a,s^{\prime})$.
+In the second equality, we explicitly formulate the operator $\tau^{\pi}$.
+In the third equality, we employ the proposition $\sum\limits_{s,a,s^{\prime}}\rho_{T}^{\pi}(s,a,s^{\prime})=1$.
+In the fourth equality, we express $\rho_{T}^{\pi}(s,a,s^{\prime})$ in the discounted sum of $\beta_{t}(s,a,s^{\prime})$.
+In the fifth equality, we employ the proposition $\beta_{t}(s,a,s^{\prime})=\beta_{t}(s,a,s^{\prime},a^{\prime})\pi(a^{\prime}|s^{\prime})$.
+In the sixth equality, we employ the proposition $\sum\limits_{s,a}\beta_{t}(s,a,s^{\prime},a^{\prime})=\beta(s^{\prime},a^{\prime})$.
+For the second term in objective (5), we define $\Phi(s,a,s^{\prime})=\alpha y(s,a,s^{\prime})$, then objective (5) can be written as:
+$\max\limits_{\pi}\min\limits_{Q(s,a)}(1-\gamma)\mathbb{E}_{s\sim\mu_{0},a\sim\pi}[Q(s,a)]+\alpha\mathbb{E}_{\rho_{M}^{\pi}}f_{\star}(\Phi(s,a,s^{\prime})/\alpha)$
+We have now completed the proof of Theorem 1.
+A.3. Proof of Theorem 2
+Next, we present the proof of Theorem 2.
+Theorem. 11.
+If $Q^{\star}(s,a)$ is the optimal solution of inner minimization in maximin problem (6), then it satisfies
+$Q^{\star}(s,a)=r(s,a,s^{\prime})-\alpha\log\dfrac{\rho^{\pi}_{M}}{\rho^{\pi_{\beta}}_{T}}-\alpha f^{\prime}(\dfrac{\rho^{\pi}_{T}}{\rho^{\pi}_{M}})+\mathcal{T}^{\pi}Q^{\star}(s,a)$ (7)
 Proof.
@@
 $\max\limits_{\pi}\min\limits_{y(s,a,s^{\prime})}\mathbb{E}_{\rho_{T}^{\pi}}[r(s,a,s^{\prime})-\alpha\log\dfrac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}-\alpha y(s,a,s^{\prime})]+\alpha\mathbb{E}_{\rho_{M}^{\pi}}[f_{\star}(y(s,a,s^{\prime}))]$
-Then the optimal solution of inner question is $y^{\star}(s,a,s^{\prime})=(f^{\prime}_{\star})^{-1}(\dfrac{\rho^{\pi}_{T}}{\rho^{\pi}_{M}})$.
-For any convex function, we have $(f^{\prime}_{\star})^{-1}(x)=f^{\prime}(x)$, so $y^{\star}(s,a,s^{\prime})=f^{\prime}(\dfrac{\rho^{\pi}_{T}}{\rho^{\pi}_{M}})$.
-Recalling that we employ the variable substitution:$y(s,a,s^{\prime})=[r(s,a,s^{\prime})-\alpha\log\frac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}+\tau^{\pi}Q(s,a)-Q(s,a)]/\alpha$, so we have
-$y^{\star}(s,a,s^{\prime})=\dfrac{r(s,a,s^{\prime})-\alpha\log\frac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}+\tau^{\pi}Q^{\star}(s,a)-Q^{\star}(s,a)}{\alpha}$
+Therefore, the optimality condition is satisfied when the derivative of $y(s,a,s^{\prime})$ equals zero, and a closed-form solution of $y^{\star}(s,a,s^{\prime})$ can be obtained: $y^{\star}(s,a,s^{\prime})=(f^{\prime}_{\star})^{-1}(\dfrac{\rho^{\pi}_{T}}{\rho^{\pi}_{M}})$.
+Let $f(x)$ be a strictly convex function, then we have $(f^{\prime}_{\star})^{-1}(x)=f^{\prime}(x)$, thus $y^{\star}(s,a,s^{\prime})=f^{\prime}(\dfrac{\rho^{\pi}_{T}}{\rho^{\pi}_{M}})$.
+Recalling that we employ the variable substitution: $y(s,a,s^{\prime})=[r(s,a,s^{\prime})-\alpha\log\frac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}+\mathcal{T}^{\pi}Q(s,a)-Q(s,a)]/\alpha$, so we have
+$y^{\star}(s,a,s^{\prime})=\dfrac{r(s,a,s^{\prime})-\alpha\log\frac{\rho_{M}^{\pi}}{\rho_{T}^{\pi_{\beta}}}+\mathcal{T}^{\pi}Q^{\star}(s,a)-Q^{\star}(s,a)}{\alpha}$
 Then the optimal solution $Q^{\star}(s,a)$ satisfies
-$Q^{\star}(s,a)=r(s,a,s^{\prime})-\alpha\log\dfrac{\rho^{\pi}_{M}}{\rho^{\pi_{\beta}}_{T}}-\alpha f^{\prime}(\dfrac{\rho^{\pi}_{T}}{\rho^{\pi}_{M}})+\tau^{\pi}Q^{\star}(s,a)$
+$Q^{\star}(s,a)=r(s,a,s^{\prime})-\alpha\log\dfrac{\rho^{\pi}_{M}}{\rho^{\pi_{\beta}}_{T}}-\alpha f^{\prime}(\dfrac{\rho^{\pi}_{T}}{\rho^{\pi}_{M}})+\mathcal{T}^{\pi}Q^{\star}(s,a)$
 Appendix B OOD Issues under Dynamics Model Integration
 OOD issues in offline model-based RL are intrinsically linked to the mismatch between synthetic and real data distributions, compounded in model-based settings, where model inaccuracies introduce additional sources of error.
-B.1 OOD Actions
+B.1. OOD Actions
 The dynamics model generates data that can significantly deviate from the distribution of the offline dataset, leading to OOD actions.
-This triggers negative model exploitation (Levine et al., 2020), resulting in poor policy performance.
+This triggers negative model exploitation Levine et al.
+(2020), resulting in poor policy performance.
 Aligning the distributions of synthetic and offline data can effectively mitigate these issues.
@@
 Specifically, we aim to match the distribution of $(s,a)_{M}^{\pi}$ with $(s,a)_{T}^{\pi_{\beta}}$, ensuring accurate value estimation of $(s,a)_{M}^{\pi}$.
-B.2 OOD states
+B.2. OOD states
 The performance of the learning policy during deployment is critically dependent on the alignment of transition pairs between the learned model and the real environment.
@@
 In summary, while optimizing the policy $\pi$ to maximize returns, we should constrain it in two directions: 1) aligning the synthetic data $(s,a,s^{\prime})_{M}^{\pi}$ with the offline data $(s,a,s^{\prime})_{T}^{\pi_{\beta}}$, and 2) ensuring that $\pi$ exhibits consistent behavior in both dynamics models and real environments.
-Appendix C Implementation Details
+Algorithm 1 DAMO
+1:Offline dataset $\mathcal{D}_{R}$, dynamics model $M$, actor network $\pi_{\phi}$, critic network $Q_{\theta}$, classifier $h_{\psi}$.
+2:Train the probabilistic dynamics model $M$ on $\mathcal{D}_{R}$
+3:Initialize $\mathcal{D}_{M}\leftarrow\varnothing$
+4:for each epoch do
+5:  Generate synthetic rollouts by model $M$
+6:  Add transitions to $\mathcal{D}_{M}$
+7:  Update $h_{\psi}$ via Eq. (3)
+8:  Calculate $\log(\rho_{M}^{\pi}/\rho_{T}^{\pi_{\beta}})$ via Eq. (4)
+9:  Update $Q_{\theta}$ by minimizing Eq. (2)
+10:  Update $\pi_{\phi}$ by maximizing Eq. (2)
+11:end for
+Appendix C Experiment Details
 In this section, we present the detailed implementation of DAMO.
-C.1 Selection of f-divergence
-As established in Theorem 5.3, the f-divergence must consistently exceed the KL-divergence to guarantee that the objective in Equation (2) serves as a lower bound for the standard RL objective $\mathcal{J}(\pi)$.
-Based on this, we specifically choose $f(x)=\frac{1}{3}(x-1)^{3}$, and its conjugate function is $f_{\star}(x)=\frac{2}{3}(x-1)^{\frac{3}{2}}.$ In principle, any choice of f-divergence is permissible, and we adhere to the selection made in the prior work (Luo et al., 2024b).
-Due to the tight schedule, we haven’t tried other choices of f-divergence.
+C.1. Algorithm of DAMO
+The complete algorithm of DAMO is summarized in Algorithm 1.
+Two replay buffers, $\mathcal{D}_{R}$ and $\mathcal{D}_{M}$, are used to store transitions from the offline dataset and model rollouts, respectively.
+C.2. Selection of f-divergence
+As established in Theorem 3, the f-divergence should consistently exceed the KL-divergence to guarantee that the objective in Equation (2) serves as a lower bound for the standard RL objective $\mathcal{J}(\pi)$.
+We adhere to the selection made in the prior work Luo et al.
+(2024a) and specifically choose $f(x)=\frac{1}{3}(x-1)^{3}$, its conjugate function is $f_{\star}(x)=\frac{2}{3}(x-1)^{\frac{3}{2}}.$ While for most values, $f(x)$ exceeds $x\log x$, in a small interval around $x=1$, $f(x)$ fails to statisfy the condition.
+In principle, it is feasible to choose a f-divergence satisfying the given condition, but due to the tight schedule, we haven’t tried other choices of f-divergence.
 However, we suppose that the choice of f-divergence won’t enormously affect the effectiveness of DAMO, since f-divergence primarily serves as a measure of discrepancy.
-Table 3: Hyperparameter of DAMO when optimizing policy
+Table 5. Hyperparameter of DAMO when optimizing policy
 Hyperparameter Value Description
@@
 $N_{iter}$ 3M Total gradient steps.
-C.2 Modification of actor training objective
+C.3. The inconsistent version of DAMO
+Section 6.1.2 focuses on evaluating the effect of DAMO’s consistent policy improvement objective.
+Prior offline MBRL methods primarily concentrate on the design of conservative Q-functions.
+Once a conservative Q-function is obtained, these methods typically apply the SAC framework for policy improvement, i.e.,
+$\max_{\pi}E_{s\sim D,a\sim\pi}Q^{\pi}(s,a)+\alpha\mathcal{H}(\pi).$
+where $D$ denotes the offline dataset and $\mathcal{H}(\pi)$ denotes the entropy of the policy $\pi$.
+In contrast, DAMO performs policy training directly based on the maximin formulation (2), where the policy improvement objective is consistent with the policy evaluation objective:
+$\max_{\pi}\min_{Q}(1-\gamma)E_{s\sim\mu_{0},a\sim\pi}[Q(s,a)]+\alpha E_{\rho_{M}^{\pi}}[f_{\star}(\Phi(s,a,s^{\prime})/\alpha)].$
+We refer to the variant of DAMO that applies the SAC-style policy improvement objective as the “inconsistent version.”
+C.4. Modification of actor training objective
 When optimizing the actor network using objective (2), we observed that the substantial residual of the Bellman error often disrupts the optimization of the actor and leads to training instability.
-To stabilize the training process, we followed the implementation of previous DICE work (Mao et al., 2024a; Sikchi et al., 2023) and fixed the hyperparameter $\alpha$ to $1.0$ during the optimization of the actor network.
+To stabilize the training process, we followed the implementation of previous DICE work Mao et al.
+(2024b); Sikchi et al.
+(2023) and fixed the hyperparameter $\alpha$ to $1.0$ during the optimization of the actor network.
 The specific form of the actor training objective is as follows:
-$\max\limits_{\pi}(1-\gamma)\mathrm{E}_{s\sim\mu_{0},a\sim\pi}[\tilde{Q}(s,a)]\\ +\mathrm{E}_{\rho_{M}^{\pi}}[f_{\star}(\Phi(s,a,s^{\prime}))]$
+$\max\limits_{\pi}(1-\gamma)\mathbb{E}_{s\sim\mu_{0},a\sim\pi}[Q(s,a)]\\ +\mathbb{E}_{\rho_{M}^{\pi}}[f_{\star}(\Phi(s,a,s^{\prime}))]$
 It should be noted that the effectiveness of DAMO does not rely on this trick. We demonstrated this in the experimental section.6.4.
-C.3 Model, classifier and Policy Optimization
+C.5. The concrete setting of halcheetah-jump
+Data collection. A SAC agent is trained in the halfcheetah task online for 1M environment steps. The reward function used in this phase is
+$r(s,a)=\max\{v_{x},3\}-0.1\times\|a\|_{2}^{2},$
+where $v_{x}$ denotes the velocity along the x-axis.
+Reward shaping. After collecting the offline dataset, we relabel the reward function to:
+$r(s,a)=\max\{v_{x},3\}-0.1\times\|a\|_{2}^{2}+15\times(z-initz),$
+where $z$ denotes the z-position of the half-cheetah and init z denotes the initial z-position.
+As the observation of the original halfcheetah environment does not contain the information of init z, we add an extra dimension to the observation to inform the agent of init z.
+Baselines.
+Three typical offline behavior-regularization methods are selected for comparison.
+BEAR Kumar et al.
+(2019)aims to constrain the policy’s actions to lie in the support of the behavioral distribution.
+BRAC Wu et al.
+(2019) is a family of algorithms that operate by penalizing the value function by some measure of discrepancy (KL divergence or MMD) between $\pi(\cdot|s)$ and $\pi^{B}(\cdot|s)$.
+TD3+BC Fujimoto and Gu (2021) introduces a behavior cloning regularization into the TD3 objective.
+C.6. Model, classifier and Policy Optimization
 For dynamics model $M$, it was depicted as a probabilistic neural network that generates a Gaussian distribution for the subsequent state and reward, contingent upon the current state and action:
 $m_{\theta}(s_{t+1},r_{t}|s_{t},a_{t})=\mathcal{N}(\mu_{\theta}(s_{t},a_{t}),\Sigma_{\theta}(s_{t},a_{t})).$
-Our model training approach is consistent with the methodology used in prior works (Yu et al., 2020; Sun et al., 2023).
+Our model training approach is consistent with the methodology used in prior works Yu et al.
+(2020); Sun et al.
+(2023).
 We train an ensemble of seven dynamics models and select the best five based on their validation prediction error from a held-out set containing 1000 transitions in the offline dataset $D_{R}$.
@@
 For classifier $h_{\psi}$, we implement the following approach: At each gradient step, we randomly sample data from both offline dataset $D_{R}$ and synthetic dataset $D_{M}$ with a fixed ratio $d$.
-This method ensures a balanced utilization of data from both DL and DG during training.
+This method ensures a balanced utilization of data from both $D_{R}$ and $D_{M}$ during training.
 Policy optimization is conducted using the Actor-Critic framework in our approach, with hyperparameter configurations aligned with previous offline model-based settings.
@@
 The specific hyperparameter settings applied for the D4RL benchmark are detailed in Table.5
-Appendix D Experiment Setting
-D.1 Benchamark
-We evaluate our approach using the D4RL benchmark (Fu et al., 2020), focusing on the “v2” version of Gym tasks.
+Appendix D Additional Experiments
+D.1. A quantitative analysis of Fig 1
+When drawing Fig 1, we applied T-SNE to perform dimensionality reduction on the data and visualized the processed data to provide an intuitive understanding of data similarity.
+During the T-SNE process, we did not quantitatively measure the non-overlapping regions between datasets.
+To enable a more rigorous comparison, we computed the Wasserstein distance between the datasets.
+The results are in the Table 6.
+Experimental results indicate that all three methods achieve a good alignment between model data and offline data, but only DAMO successfully aligns model data with real data, which is consistent with our analysis in the paper.
+Comparing the directly computed Wasserstein distance with the Wasserstein distance after T-SNE dimensionality reduction, we observe that although the absolute values change, the relative ordering remains the same, confirming the validity of using T-SNE for visualization.
+Table 6.
+Wasserstein distance between the offline, real, and model data.
+The first column clarifies the type of data, for example, Offline-Model means the Wasserstein distance between Offline data and Model data.
+’w/ Tsne’ represents the usage of T-SNE to reduce the dimension of data and then calculate the Wasserstein distance.
+Distance MOPO MOBILE DAMO (ours)
+Offline-Model 0.24 0.13 0.16
+Model-Real 0.39 0.43 0.27
+Offline-Model(w/ T-SNE) 2.17 1.96 2.01
+Model-Real(w/ T-SNE) 10.28 6.76 4.35
+D.2. The performance in noisy environments
+Table 7. Results on the noisy mujoco.
+Task Name MOPO MOBILE DAMO (ours)
+Halfcheetah-me 42.10 72.53 $74.51\pm 7.55$
+Hopper-me 63.82 93.86 $95.36\pm 4.14$
+Walker2d-me 108.69 114.90 $113.94\pm 1.15$
+Noisy environment refers to a situation where random noise interferes with the agent’s state observation during testing, which is normal in real-world applications.
+Therefore, we have supplemented the following experiment in Mujoco to test whether DAMO is robust enough to demonstrate strong performance in noisy environment.
+We still use DAMO to train the policy on clean offline data, but during the testing phase, at each step transition, the environment introduces noise to the state observed by the agent with a certain probability.
+Under this setting, we tested the performance of DAMO, MOPO, and MOBILE, and the results are in the table 7.
+Both DAMO and MOBILE exhibit robust performance in noisy environments, effectively mitigating the impact of environmental noise on policy performance, whereas MOPO experiences a significant performance degradation.
+Notably, injecting noise in the Walker2d environment does not lead to a noticeable decline in policy performance, which may be attributed to the characteristics of the environment and task, warranting further investigation.
+D.3. DAMO with different hyperparameters for dual alignment
+Using different parameters for the two regularization terms in the DAMO can serve as a trick to improve algorithm performance.
+We conducted an experiment under the halfcheetah-medium-expert dataset to evaluate its impact.
+We set the coefficient of the term $D_{f}(\rho_{T}^{\pi}\|\rho_{M}^{\pi})$ in (1) to be $\beta$ and keep the coefficient of the term $\log(\rho_{M}^{\pi}/\rho_{T}^{\pi_{\beta}})$ to be $\alpha$.
+The training process is illustrated in Fig 6.
+This trick does not yield a significant effect on the Halfcheetah-Medium-Expert dataset.
+This may be attributed to two factors: DAMO itself is insensitive to the parameters $\alpha$ and $\beta$, and the Halfcheetah environment also exhibits low sensitivity to these parameters.
+The final scores from top to bottom are 103.1, 101.78, and 100.33.
+Figure 6. fig-Setting different parameters for two penalization terms
+Appendix E Experiment Setting
+E.1. Benchamark
+In the D4RL benchmark Fu et al.
+(2020), we focus on the “v2” version of Gym tasks.
 For DICE-based methods (Algae-DICE, Opti-DICE, and O-DICE), we directly adopt the scores reported in their respective original papers.
 However, since the results for O-DICE on the random dataset are not provided in the original paper, we exclude them from our comparison.
-For offline model-based methods, we retrained MOPO (Yu et al., 2020) and MOBILE (Sun et al., 2023) on the “v2” Gym datasets.
+For offline model-based methods, we retrained MOPO Yu et al.
+(2020) and MOBILE Sun et al.
+(2023) on the “v2” Gym datasets.
 For the remaining methods, we directly report the scores provided in their original papers.
 All results are summarized in Table 1.
-Table 4: Hyperparameters of DAMO used in the D4RL datasets.
+In the NeoRL benchmark Qin et al.
+(2022), we report the performance of BC, CQL, and MOPO based on the original NeoRL paper, while the performance of TD3+BC, EDAC, and MOBILE is derived from the scores reported in the MOBILE Sun et al.
+(2023) paper.
+Table 8. Hyperparameters of DAMO used in the D4RL datasets.
 Task Name $\alpha$ $K$ $l_{r}$ of actor
@@
 walker2d-medium-expert 1.5 3 $1\times 10^{-4}$
-D.2 Hyperparameters
-We provide a list of the hyperparameters that were tuned during our experiments. The detailed configurations are summarized in Table 4.
+E.2. Hyperparameters
+We provide a list of the hyperparameters that were tuned during our experiments.
+The detailed configurations in the D4RL benchmark are summarized in Table 8.
+The detailed configurations in the NeoRL benchmark are summarized in Table 9.
+It should be noted that we considered two additional hyperparameters in the NeoRL benchmark, due to the narrow data coverage and noisy environments.
 Coefficient $\alpha$.
@@
 Rollout Length $K$. Similar to MOPO, we employ short-horizon branch rollouts in DAMO. For Gym tasks in the D4RL benchmark, the horizon parameter $K$ is tuned within the range of $\{1,3,5\}$.
-The actor learning rate $l_{r}$ Since DAMO operates as a Maximin optimization problem and is susceptible to saddle points, we optimize the actor’s learning rate to achieve robust performance.
+The actor learning rate $l_{r}$.
+Since DAMO operates as a Maximin optimization problem and is susceptible to saddle points, we optimize the actor’s learning rate to achieve robust performance.
 The choice of learning rate plays a pivotal role in DAMO’s effectiveness.
 We tune the actor learning rate $l_{r}$ in the range of $\{1\times 10^{-5},5\times 10^{-5},1\times 10^{-4}\}$.
-D.3 Computing Infrastructure
+Offline data ratio.
+The limited coverage of the transition space leads to poor performance of the trained dynamics in some datasets.
+Consequently, we have to downweight the ratio of synthethic data used to optimize the objective for stabilizing the training process.
+We tune the offline data ratio within the range of $\{0.05,0.1,0.3\}$.
+Auto-alpha. Similar to MOBILE, we forbid the automatic entropy tuning trick in some datasets.
+Table 9. Hyperparameters of DAMO used in the NeoRL datasets.
+Task Name $\alpha$ $K$ $l_{r}$ of actor Offline data ratio Auto-alpha
+Halfcheetah-L 0.3 5 $1\times 10^{-4}$ 0.05 True
+Hopper-L 1 5 $1\times 10^{-4}$ 0.05 True
+Walker2d-L 2.5 1 $1\times 10^{-4}$ 0.1 True
+Halfcheetah-M 0.5 3 $1\times 10^{-4}$ 0.05 True
+Hopper-M 3 5 $5\times 10^{-5}$ 0.3 False
+Walker2d-M 3.5 1 $1\times 10^{-5}$ 0.3 True
+Halfcheetah-H 3 3 $1\times 10^{-4}$ 0.05 True
+Hopper-H 5 5 $1\times 10^{-5}$ 0.05 False
+Walker2d-H 3 3 $1\times 10^{-4}$ 0.3 True
+E.3. Computing Infrastructure
 The experimental setup includes a GeForce GTX 3090 GPU and an Intel(R) Xeon(R) Gold 6330 CPU at 2.00GHz.